### v1

# Accelerated Ostwald ripening by chemical activity

Abstract
Phase separation of biomolecular condensates is ubiquitous in living cells, promoting colocalization of enzymes and their substrates as well as achieving membrane-free compartmentalization. Energy-consuming processes are routinely used to regulate biocondensate growth by opposing the thermodynamic tendency toward coarsening. At the same time, cells often use energy to instead accelerate thermodynamic processes. Here, we theoretically explore the possibility of utilizing chemical reactions to accelerate biocondensate coarsening. We combine Lifshitz-Slyozov theory with a reaction-diffusion approach, wherein particles interconvert between phase-separating and inert forms. We find that mass conservation restricts the volume growth to be linear in time (as in the passive case) despite activity, though if reactions are restricted to occur only outside droplets, the rate of Ostwald ripening can be increased by an arbitrarily large factor. Our acceleration theory is quantitatively supported by recent experiments on ripening in the presence of fueled interconversion reactions, under precisely the predicted conditions. We posit that the ability to induce rapid biocondensate coarsening can be advantageous in synthetic-biological contexts as a regulator of metabolic channeling.
I Introduction
Ostwald ripening refers to the thermodynamic process of coarsening of solute-rich phase-separated droplets. As a result of surface tension, smaller droplets shrink due to their larger Laplace pressure whereas larger droplet grow via the diffusion of solute. Average droplet size may also increase due to coalescence events. These processes continue in principle until phase-separation completes. However, a vastly richer physics emerges when activity and energy consumption are introduced. One well-studied outcome is “size control” [1, 2, 3, 4, 5], whereby active processes prevent unbounded coarsening. Size control can be achieved via nonequilibrium reactions that interconvert the solute between phase-separating and inert forms. Additional possible outcomes of activity include arrested ripening [6] (where small droplets do not shrink), creation of Turing patterns [7, 8] (where the dense phase is destabilized by the coupling between thermodynamics and the reactions), and reversal of Ostwald ripening [9] (where dilute bubbles form inside the dense phase). Further complexities such as nonlocal elasticity [10, 11, 12, 13] or charge asymmetries [14] extend the list of possible outcomes. In contrast to these effects which oppose Ostwald ripening, it was recently demonstrated experimentally that fueled interconversion reactions in organochemical emulsions can subtantially accelerate ripening [15]. How do these interconversion reactions lead to such a different outcome?
One motivation to investigate accelerated Ostwald ripening is its potential biological relevance. Phase separation of biomolecules to form biomolecular condensates (also termed “membraneless organelles”) is pervasive in cells, enabling the compartmentalization of biochemical processes [16, 17]. Biocondensate size can be crucial for function [18, 19, 20, 21, 22]. For example, consider the algal pyrenoid [23] — a condensate composed of Rubisco (a carbon-fixating enzyme) and EPYC1 (a “sticky” linker protein that drives phase separation) — which is dynamically regulated through energy-consuming phosphorylation cycles. During cell division, phosphorylation renders EPYC1 less sticky, breaking the pyrenoid into small droplets via a size-control mechanism, thus favoring equipartition of Rubisco between the daughter cells. After division, dephosphorylation allows recoarsening into a single pyrenoid in each daughter cell. For cases like the pyrenoid where a single condensate, or a small number of large condensates, is advantageous, might cells use energy to accelerate Ostwald ripening, and how might this be achieved?
Figure 1: Illustration of the model. (a) Passive system: A slightly supersaturated, monodisperse suspension of “sticky” particles. Spontaneously-formed dense-phase nuclei will coarsen at the expense of the material available in the surrounding dilute phase. Once the supersaturation becomes small, Ostwald ripening commences — while larger droplets keep growing, smaller droplets, whose Laplace pressure is higher, begin shrinking. (b) Active system: Motivated by the active ripening of the algal pyrenoid , we introduce kinases and phosphatases to the above suspension. The former phosphorylate the sticky proteins, rendering them “nonsticky” (inert), and the latter dephosphorylate nonsticky proteins returning them to the sticky form. In our study, we investigate whether these active processes can allow for faster Ostwald ripening.
To answer this question, we consider a solution consisting of a “sticky” species that tends to phase separate and an inert “nonsticky” species, into which we introduce interconversion reactions. We devise a general theory which covers all possible outcomes resulting from these reactions, and identifies the only regime in which it is possible to achieve acceleration. In general, we find that reactions in the dilute phase promote the influx of sticky particles, whereas conversions inside the dense droplets oppose droplet growth. In line with the experiments in Ref. [15], we find that when the reactions inside the droplet are negligible, an acceleration by an arbitrary factor is possible. However, a key prediction is that a time scaling faster than Lifshitz-Slyozov scaling [24] is physically unrealizable, and only the coefficient can be tuned by activity. Our theory, among others, also yields the previously reported phenomena of size control [3, 4, 5], ripening arrest [6], and droplet instabilities [7, 8, 9].
In Sec. II, we recall the standard theoretical toolkit for studying passive Ostwald ripening in a binary mixture, including thermodynamics of coexistence, linear-irreversible thermodynamics, and a set of approximations representing diffusion-limited growth. In Sec. III, we extend the theory to a ternary mixture which undergoes chemical reaction. In Sec. IV we solve the preceding equations, specify the dimensionless parameters separating the possible outcomes, and explore the behavior in the unique regime of acceleration along with a comparison to experiment [15]. Finally, in Sec. V we summarize our results, note the limitations of our theory, discuss its applicability to biology and put it in a broader theoretical context. All technical calculations that are not shown in these sections can be found in SI Appendix.
II Passive theory
In this section, we lay out the general theoretical foundation for studying Ostwald ripening. We will consider a binary fluid — only sticky molecules surrounded by an inert solvent — in the absence of an additional nonsticky component and reactions (see illustration in Fig. 1(a)). The aim of this section is to recall the classical Lifshitz-Slyozov ripening dynamics [24]. In the subsequent Secs. III and IV, we investigate whether introducing sticky-nonsticky interconversion reactions could allow us to accelerate these dynamics; cf. Fig. 1(b).
We are interested in the intermediate stages of phase separation: The droplets have already nucleated and have been coarsening for some time, so they are large (compared to the dense-dilute interface width), spherical, and sparse. Surface tension leads to ripening — larger droplets grow and smaller ones shrink by diffusion-driven exchange of material with the surrounding dilute phase. (We ignore droplet diffusion and coalescence [25, 26].) We present the standard linear-irreversible thermodynamics approach for modeling the dynamics of the particle concentration field in the vicinity of a droplet, from which we then obtain the droplet radius dynamics [24, 3, 1].
II.1 Setup
We begin by enumerating the components of the system and the governing thermodynamic principles. We consider a solution of “sticky” particles (e.g., proteins), labeled $\mathrm{s}$, which tend to phase separate. These particles are immersed in an inert liquid (e.g., the cytoplasm) whose volume $V$ and temperature $T$ are fixed. We denote the local volume fraction of the sticky species as $\phi_{\mathrm{s}}$. We assume that the entire system is incompressible, such that the local solvent volume fraction is $1-\phi_{\mathrm{s}}$. [^1] We assume that the cytoplasm has a fixed total concentration of macromolecules immersed in water. Of this concentration, $\phi_{\mathrm{s}}$ is the volume fraction of the sticky protein of interest, and $1-\phi_{\mathrm{s}}$ is thus the volume fraction of the remaining “irrelevant” (indistinguishable and inert) macromolecules. The equilibrium state of the system is, therefore, uniquely determined by a normalized free energy per molecule, $f=F/(Nk_{\mathrm{B}}T)$, which is a function of the sticky volume fraction, $f(\phi_{\mathrm{s}})$ [27].
Phase separation is possible whenever the free energy has more than one stable volume fraction. We denote the coexistence volume fraction of the dense phase (sticky-rich, “liquid”) by $\phi_{\mathrm{s}\mathrm{l}}^{\circ}$ and of the dilute phase (sticky-poor, “gas”) by $\phi_{\mathrm{s}\mathrm{g}}^{\circ}$. When two phases in contact are in equilibrium, the chemical potential and pressure of the two phases equate [27].
Since Ostwald ripening is driven by surface tension, we distinguish between the above “bulk” coexistence volume fractions, $\{\phi_{\mathrm{s}\mathrm{l}}^{\circ},\phi_{\mathrm{s}\mathrm{g}}^{\circ}\}$, and the volume fractions found around droplets of finite radius $R$, $\{\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}(R),\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)\}$. These volume fractions differ since the liquid pressure must now oppose both the gas pressure and so-called Laplace pressure, the latter associated with droplet curvature [28]. In the limit of large droplets compared to the interface width, the two gas-phase volume fractions are generically connected via $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)=\phi_{\mathrm{s}\mathrm{g}}^{\circ}+\lambda/R$, where $\lambda$ is a Laplace-pressure correction term, with a similar expression for the liquid-phase volume fraction. In the SI Appendix, Sec. I, we recall that $\lambda$ is proportional to the surface tension and the compressibility of each phase [29, 1]. In the strong phase separation limit, one can expect $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}\simeq 1$ and $\phi_{\mathrm{s}\mathrm{g}}^{\circ}\lesssim\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}\ll 1$; see SI Appendix, Sec. I. To enable phase separation, the system is prepared in some initial supersaturated volume fraction denoted $\varphi^{\infty}$.
II.2 Linear-irreversible thermodynamics
We proceed to extend the above thermodynamic treatment and connect it to the dynamics of the droplets. To address the fact that a collection of dense droplets is formed within a dilute medium (cf. Fig. 1(a)), we adopt a coarse-grained view where a given particle configuration is modeled via a continuous, nonuniform volume fraction field. Following density functional theory, we approximate the free energy as a functional of the volume fraction field [30, 31, 2],
$F=k_{\mathrm{B}}T\int\mathrm{d}\mathbf{r}\left[f(\phi_{\mathrm{s}}(\mathbf{r}))+\frac{\kappa_{\mathrm{s}}}{2}|\boldsymbol{\nabla}\phi_{\mathrm{s}}(\mathbf{r})|^{2}\right],$ (1)
where we have included a leading-order, quadratic energy penalty for gradients in the sticky species. The surface tension and interface width are related to the prefactor $\kappa_{\mathrm{s}}>0$ [29]; see SI Appendix, Sec. I. The volume fraction fields therefore modulate smoothly from $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}$ in dense-phase regions to $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$ in dilute regions.
To study droplet dynamics, we follow the lines of linear-irreversible thermodynamics, whereby the conserved volume fraction fields, $\partial\phi_{\mathrm{s}}/\partial t=-\boldsymbol{\nabla}\cdot\mathbf{J}_{\mathrm{s}}$, respond linearly to local gradients in the chemical potentials governed by the free energy functional [32, 33], $\mathbf{J}_{\mathrm{s}}=-\Gamma_{\mathrm{s}\mathrm{s}}\boldsymbol{\nabla}\delta F/\delta\phi_{\mathrm{s}}$, where $\mathbf{J}_{\mathrm{s}}$ is the local sticky flux vector, $\Gamma_{\mathrm{s}\mathrm{s}}$ is the mobility coefficient, and $\delta F/\delta\phi_{\mathrm{s}}$ is the local sticky chemical potential. Explicitly [1],
$\frac{\partial\phi_{\mathrm{s}}(\mathbf{r},t)}{\partial t}$ $=$ $\boldsymbol{\nabla}\cdot\Biggl{\{}\Gamma_{\mathrm{s}\mathrm{s}}(\phi_{\mathrm{s}}(\mathbf{r},t))\boldsymbol{\nabla}\Biggl{[}\frac{\partial f}{\partial\phi_{\mathrm{s}}}\Biggr{|}_{\phi_{\mathrm{s}}(\mathbf{r},t)}$ (2)
$-\kappa_{\mathrm{s}}\nabla^{2}\phi_{\mathrm{s}}(\mathbf{r},t)\Biggr{]}\Biggr{\}}.$
II.3 Quasistatic growth
Here, we introduce a number of standard simplifications that allow for analytical treatment of late-stage coarsening [31, 2, 1]. First, we assume that droplets are sparse; ripening occurs through diffusive exchange of material with a slightly supersaturated dilute phase, $0<\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}\ll 1$. Owing to the sparsity, in what follows, we may concentrate on a single droplet, ignoring all others.
We approximate the droplet to be spherical owing to surface tension, with radius $R(t)$ much larger than the droplet interface. In the vicinity of the droplet, we thus split the volume fraction fields and fluxes into interior (dense-phase) and exterior (dilute-phase) profiles, $\phi_{\mathrm{s}\mathrm{l}}(r,t)=\phi_{\mathrm{s}}(|\mathbf{r}-\mathbf{r}_{\mathrm{CM}}|<R(t),t)$ and $\phi_{\mathrm{s}\mathrm{g}}(r,t)=\phi_{\mathrm{s}}(|\mathbf{r}-\mathbf{r}_{\mathrm{CM}}|>R(t),t)$, where $\mathbf{r}_{\mathrm{CM}}$ is the coordinate of the droplet center.
To stich the two profiles via the appropriate boundary conditions, we assume that the surface thermodynamics is “fast”. Namely, growth of large droplets requires diffusive rearrangement of material over lengthscales $\sim R$, whereas the response in the narrow interface involves smaller distances $\sim\kappa_{\mathrm{s}}^{1/2}$. Therefore, the interfaces surrounding droplets are thermodynamically equilibrated at every instant during growth (i.e., we ignore “interface resistance” [34, 35]). We thus require $\phi_{\mathrm{s}\mathrm{l}}(R(t),t)=\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}(R(t))$, and $\phi_{\mathrm{s}\mathrm{g}}(R(t),t)=\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R(t))$ for every instantaneous radius $R(t)$. Further, far away from a given droplet, we impose the instantaneous saturation, $\phi_{\mathrm{s}\mathrm{g}}(\infty,t)=\varphi^{\infty}$.
Lastly, droplet growth requires diffusive rearrangement of sticky volume $\sim(\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}})\sim 1$, whereas the surrounding gas phase involves the much smaller difference $\sim(\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}})\ll 1$. Therefore, we approximate the droplet growth dynamics to be quasistatic: The volume fraction profiles $\{\phi_{\mathrm{s}\mathrm{l}}(r;R),\phi_{\mathrm{s}\mathrm{g}}(r;R)\}$ are first found for a quasi-steady state of fixed $R$ using the above boundary conditions. Next, the radius growth $\mathrm{d}R/\mathrm{d}t$ is determined by the quasistatic volume-fraction gradient corresponding to that radius. Since the supersaturation is close to $\phi_{\mathrm{s}\mathrm{g}}^{\circ}$, we linearize Eq. (2) around the binodal in each phase, $\phi_{\mathrm{s}\alpha}^{\mathrm{L}}(R)$ ($\alpha=\mathrm{l},\mathrm{g}$).
Upon linearizing Eq. (2) and taking the quasistatic and spherically-symmetric limit, the volume fractions $\phi_{\mathrm{s}\mathrm{l}}(r)$ and $\phi_{\mathrm{s}\mathrm{g}}(r)$ simply obey the stationary diffusion equation,
$\frac{D_{\alpha}}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}\frac{\partial\phi_{\mathrm{s}\alpha}}{\partial r}\right)=0,$ (3)
with the above boundary conditions, where for transparency we kept the diffusion constant
$D_{\alpha}=\left.\Gamma_{\mathrm{s}\mathrm{s}}\frac{\partial^{2}f}{\partial\phi_{\mathrm{s}}\partial\phi_{\mathrm{s}}}\right|_{\phi_{\mathrm{s}\alpha}^{\mathrm{L}}}$ (4)
which emerged from the linearization, and we neglected the surface tension (bilaplacian) term away from the interface due to the smooth gradients ($\kappa_{\mathrm{s}}/R^{4}\ll 1/R^{2}$). With these approximations, a droplet evolves in time as
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{J_{\mathrm{s}\mathrm{l}}(R(t))-J_{\mathrm{s}\mathrm{g}}(R(t))}{\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}(R(t))-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R(t))},$ (5)
where $J_{\mathrm{s}\alpha}(R)=D_{\alpha}(\partial\phi_{\mathrm{s}\alpha}/\partial r)|_{r=R}$ are the outward fluxes in the radial direction within the $\alpha$th phase at the interface, which we will compute in the upcoming section. Eq. (5) can be understood by a material conservation argument: $J_{\mathrm{s}\mathrm{l}}-J_{\mathrm{s}\mathrm{g}}$ is the net amount of sticky volume which is being brought further away from the droplet’s center, per unit time, per unit area. So to convert this material displacement to length, this net flux is divided by the difference between the dense and dilute sticky volume fractions, $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$. For a derivation, see SI Appendix, Sec. II.
II.4 Droplet dynamics
In this section, we find the quasistatic volume fraction field, and obtain explicit expressions for the droplet radius dynamics. Upon introducing material conservation, we remind why ripening occurs instead of simple growth, and cite the Lifshitz-Slyozov law.
With the above boundary conditions, we find the solution to Eq. (3),
$\phi_{\mathrm{s}\mathrm{l}}(r;R)$ $=$ $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}(R),$ (6)
$\phi_{\mathrm{s}\mathrm{g}}(r;R)$ $=$ $\varphi^{\infty}-[\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)]\frac{R}{r},$ (7)
and the resultant diffusive gas-phase flux into the droplet,
$J_{\mathrm{s}\mathrm{g}}(R)=-\frac{D_{\mathrm{g}}}{R}\left[\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)\right],$ (8)
whereas $J_{\mathrm{s}\mathrm{l}}(R)=0$. Finally, recalling the leading-order partition coefficient $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}(R)-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)\simeq 1$, we obtain the radius growth dynamics from Eq. (5[24]
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R}\left[(\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})-\frac{\lambda}{R}\right].$ (9)
Thus, large droplets ($R>\lambda/(\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})$) grow ($\mathrm{d}R/\mathrm{d}t>0$), while small droplets ($R<\lambda/(\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})$) are thermodynamically unfavorable due to excessive Laplace pressure, and therefore shrink ($\mathrm{d}R/\mathrm{d}t<0$).
Note that the total amount of sticky material is conserved, thus implicitly connecting the saturation $\varphi^{\infty}$ to the radii. Namely, defining the droplet-radius distribution $f(R,t)$ (normalized such that $V\int_{0}^{\infty}\mathrm{d}Rf(R,t)$ is the total number of droplets in the container), material conservation across the sparse droplets and their dilute surroundings can be approximately written as [24]
$\varphi^{\infty}(t)+\phi_{\mathrm{s}\mathrm{l}}^{\circ}\int_{0}^{\infty}\mathrm{d}Rf(R,t)\frac{4}{3}\pi R^{3}=\mathrm{const}.$ (10)
As droplets grow, they take up sticky particles from the supersaturated gas phase, meaning that the saturation $\varphi^{\infty}$ decreases over time. Therefore, according to Eq. (9), droplets that were previously able to grow may become unstable to Laplace pressure and will now shrink, while the larger droplets continue growing. (This process continues until a single large droplet is found such that $\varphi^{\infty}=\phi_{\mathrm{s}\mathrm{g}}^{\circ}$.) Thus, the Lifshitz-Slyozov scaling law emerges [24]: The overall number of droplets decreases in time (as the smaller droplets vanish), the saturation decreases too, while the average droplet radius grows as
$\langle R(t)\rangle^{3}-\langle R(0)\rangle^{3}=\frac{4}{9}D_{\mathrm{g}}\lambda t.$ (11)
We now investigate whether chemical reactions may accelerate the ripening law of Eq. (11).
III Active theory
We introduce chemical interconversion reactions as follows. We again assume that a supersaturated solution containing a density $\varphi^{\infty}$ of sticky particles is prepared, as previously; cf. Fig. 1(a). To it, we add an enzyme (e.g., a kinase) that converts sticky proteins into nonsticky ones, and vice versa by another enzyme (e.g., a phosphatase); cf. Fig. 1(b). We proceed to extend the passive theory to include these chemical reactions and the subsequent presence of the third nonsticky component. The active theory constructed below will be compared to the well-known growth laws summarized above.
We define the total reaction rate, $s(\{\phi_{i}\})$, as a function of the local particle concentrations ($i=\mathrm{s},\mathrm{ns}$), with associated one-to-one stoichiometric coefficients $\upsilon_{\mathrm{s}}=-1$ and $\upsilon_{\mathrm{ns}}=1$. In SI Appendix, Sec. II, we generalize the type of reactions one may consider.
III.1 Setup
As nonsticky proteins appear, the normalized free energy per molecule should depend on both sticky and nonsticky volume fractions, $f(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})$. By nonsticky we imply that these particles are inert, meaning that phase separation is driven only by the sticky species. We denote the coexistence volume fractions of the dense phase (sticky-rich, “liquid”) by $\{\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\}$ and of the dilute phase (sticky-poor, “gas”) by $\{\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\}$. When the two phases in contact are in equilibrium, we now require equal sticky chemical potentials, nonsticky chemical potentials, and pressures. In accordance with the Gibbs phase rule [36], these three conditions are insufficient to determine all four volume fractions. Therefore, one should also impose material conservation: With a known total amount of each species, there are two more constraints, sufficient to determine all four volume fractions as well as the equilibrium fraction of the dense phase.
However, during coarsening, the volume fractions are not uniform everywhere, so a local dynamical constraint must be imposed around each droplet instead. Namely, as a droplet grows, the net amount of sticky material joining the droplet, as captured by Eq. (5), must track with the net amount of nonsticky material being expelled [37],
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{J_{\mathrm{ns}\mathrm{l}}(R(t))-J_{\mathrm{ns}\mathrm{g}}(R(t))}{\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}(R(t))-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R(t))},$ (12)
where we will compute the fluxes $J_{i\alpha}(R)$ ($i=\mathrm{s},\mathrm{ns}$ and $\alpha=\mathrm{l},\mathrm{g}$) in Sec. III.4. Eq. (5) and Eq. (12), together with the thermodynamic constraints, are sufficient to determine all four volume fractions as well as the sought-after focal droplet growth rate. In the strong phase separation limit of sticky species, the volume fractions $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}\simeq 1$ and $\phi_{\mathrm{s}\mathrm{g}}^{\circ}\lesssim\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}\ll 1$ remain unchanged, with $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\simeq 0$ as a result, and so only $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\ll 1$ is unknown; see SI Appendix, Sec. I.
III.2 Reaction-diffusion approach
Combined with the linear-irreversible thermodynamics description of Sec. II.2, the reaction-diffusion dynamics we investigate henceforth is (with $i,j,k=\mathrm{s},\mathrm{ns}$)
$\frac{\partial\phi_{i}(\mathbf{r},t)}{\partial t}$ $=$ $\boldsymbol{\nabla}\cdot\Biggl{\{}\sum_{j}\Gamma_{ij}(\{\phi_{k}(\mathbf{r},t)\})\boldsymbol{\nabla}\Biggl{[}\frac{\partial f}{\partial\phi_{j}}\Biggr{|}_{\{\phi_{k}(\mathbf{r},t)\}}$ (13)
$-\kappa_{j}\nabla^{2}\phi_{j}(\mathbf{r},t)\Biggr{]}\Biggr{\}}+\upsilon_{i}s(\{\phi_{k}(\mathbf{r},t)\}),$
where the first term is the conservative flux in linear response to local chemical potential gradients, arising from the free energy functional [^2] We omitted energetic cost in $\boldsymbol{\nabla}\rho_{\mathrm{ns}}(\mathbf{r})$ as the nonsticky proteins are inert.
$F=k_{\mathrm{B}}T\int\mathrm{d}\mathbf{r}\left[f(\phi_{\mathrm{s}}(\mathbf{r}),\phi_{\mathrm{ns}}(\mathbf{r}))+\frac{\kappa_{\mathrm{s}}}{2}|\boldsymbol{\nabla}\phi_{\mathrm{s}}(\mathbf{r})|^{2}\right],$ (14)
and the second term describes reactions. Note that $\phi_{\mathrm{s}}(\mathbf{r},t)+\phi_{\mathrm{ns}}(\mathbf{r},t)$ is unaffected by the reactions and is thus a conserved field.
We assume that the switching rates are sufficiently slow such that the reactions do not affect the surface tension [38, 39]. In other words, the free-energy-driven phase separation continues unchanged, and the reactions only modulate the diffusive fluxes via which droplet growth or shrinkage occurs.
III.3 Quasistatic growth
Here, we build upon the assumptions of Sec. II.3 to enable analytical treatment of the reaction-diffusion framework considered here [2, 1].
As before, we assume that the droplets are sparse, so we still concentrate on a focal droplet which grows or shrinks by diffusion of sticky and nonsticky material. The saturation $\varphi^{\infty}$, however, is now divided between sticky and nonsticky particles, $\varphi^{\infty}_{\mathrm{s}}$ and $\varphi^{\infty}_{\mathrm{ns}}$, respectively, such that $\varphi_{\mathrm{s}}^{\infty}+\varphi_{\mathrm{ns}}^{\infty}=\varphi^{\infty}$. Far away from the droplet, the reactions equilibrate, so $s(\varphi_{\mathrm{s}}^{\infty},\varphi_{\mathrm{ns}}^{\infty})=0$.
We once again assume that droplets are spherical, with radius $R(t)$ much larger than the interface width. Thus, we split the volume fraction fields into $\phi_{i\mathrm{l}}(r,t)=\phi_{i}(|\mathbf{r}-\mathbf{r}_{\mathrm{CM}}|<R(t),t)$ and $\phi_{i\mathrm{g}}(r,t)=\phi_{i}(|\mathbf{r}-\mathbf{r}_{\mathrm{CM}}|>R(t),t)$ for each species $i=\mathrm{s},\mathrm{ns}$.
As mentioned in Sec. III.2, we suppose that the reactions are not rapid enough to interfere with the surface thermodynamics [38, 39]. Namely, the reactive lengthscale $\ell$, representing the distance that a particle diffuses in between switching reactions, is assumed much larger than $\kappa_{\mathrm{s}}^{1/2}$, so the fast-surface-thermodynamics assumption of Sec. II.3 carries over. We thus require $\phi_{i\alpha}(R(t),t)=\phi_{i\alpha}^{\mathrm{L}}(R(t))$ ($i=\mathrm{s},\mathrm{ns}$ and $\alpha=\mathrm{l},\mathrm{g}$) for every instantaneous radius $R(t)$.
Lastly, we again approximate the droplet growth to be quasistatic: The volume fraction profiles $\{\phi_{i\alpha}(r;R)\}$ are found for a quasi-steady state of fixed $R$. In turn, the radius growth $\mathrm{d}R/\mathrm{d}t$ is determined by Eqs. (5) and (12) where the volume fractions and flux are evaluated for the current $R(t)$. At this point, in Sec. II.3 we proceeded to linearize the diffusion equation around $\{\phi_{i\alpha}^{\mathrm{L}}(R)\}$ for each phase and species. Below, we will show the outline for this procedure in the gas phase. For the liquid phase, however, the reactions render the linearization much more intricate, and we defer details to SI Appendix, Sec. II.
In the gas phase, we assume that the reduced $\varphi^{\infty}_{\mathrm{s}}$ is close to the binodal $\phi_{\mathrm{s}\mathrm{g}}^{\circ}$. Therefore, we linearize Eq. (13) around the coexistence volume fractions $\phi_{i\mathrm{g}}(r;R)=\phi_{i\mathrm{g}}^{\mathrm{L}}(R)+\delta\phi_{i\mathrm{g}}(r;R)$ ($i=\mathrm{s},\mathrm{ns}$). Since the gas is dilute, the sticky and nonsticky species are sparse, effectively noninteracting, and similar in size, so the diffusion matrix can be approximated as a unit matrix, $\sum_{k}\Gamma_{ik}(\partial^{2}f/\partial\phi_{k}\partial\phi_{j})|_{\{\phi_{l\alpha}^{\mathrm{L}}\}}=D_{\mathrm{g}}\delta_{ij}$, with $D_{\mathrm{g}}$ given by Eq. (4). Furthermore, since the concentrations in the gas phase are small, and no reactions occur in the absence of material ($s(0,0)=0$), we keep only the first-order contribution to the conversion reactions, $s(\phi_{\mathrm{s}\mathrm{g}},\phi_{\mathrm{ns}\mathrm{g}})\simeq k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}$, where $k_{\mathrm{s}\to\mathrm{ns}},k_{\mathrm{ns}\to\mathrm{s}}>0$ are effective rate constants. [^3] With our choice of stoichiometry, $\upsilon_{\mathrm{s}}=-1$ and $\upsilon_{\mathrm{ns}}=+1$, this choice of signs for $k_{i\to j}$ is required for the chemical reactions’ equilibrium to be stable; see details in SI Appendix, Sec. II. With this, the linearized equations become (SI Appendix, Sec. II)
$\frac{D_{\mathrm{g}}}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}\frac{\partial\phi_{j\mathrm{g}}}{\partial r}\right)=-\upsilon_{i}\left(k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}}\right).$ (15)
Note that for the above reaction rule, the saturations read
[…]
Compared to the passive scenario of Sec. II.3 with only sticky particles, the saturation of the sticky species has decreased from $\varphi^{\infty}$ to $\varphi_{\mathrm{s}}^{\infty}$. Does this loss of saturation still allow for accelerated Ostwald ripening?
III.4 Droplet dynamics
In this section, we find the quasistatic volume-fraction fields, and obtain the relevant expressions for the focal-droplet radius dynamics under the effect of reactions.
We find the solution to Eqs. (15),
$\phi_{\mathrm{s}\mathrm{g}}(r;R)+\phi_{\mathrm{ns}\mathrm{g}}(r;R)$
$\quad=\varphi^{\infty}-\left[\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)\right]\frac{R}{r},$ (17)
$k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}(r;R)-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}(r;R)$
$\quad=\left[k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)\right]\frac{R}{r}e^{-(r-R)/\ell}.$ (18)
Eq. (17) is reminiscent of Eq. (7), as the total volume fraction of proteins is not affected by the interconversion reaction, and hence its profile is simply diffusive as in the passive case. In Eq. (18), we identified explicitly the reactive lengthscale — the distance that a particle diffuses in between reactions [3],
$\ell=\sqrt{\frac{D_{\mathrm{g}}}{k_{\mathrm{ns}\to\mathrm{s}}+k_{\mathrm{s}\to\mathrm{ns}}}}.$ (19)
It sets the distance over which the reactions equilibrate ($s(\phi_{\mathrm{s}\mathrm{g}}(r;R),\phi_{\mathrm{ns}\mathrm{g}}(r;R))=0$). As mentioned above, for the assumption of fast surface thermodynamics to hold, we need $\ell\gg\kappa_{\mathrm{s}}^{1/2}$. In Sec. IV we will investigate two regimes, $R\ll\ell$ and $R\gg\ell$. Upon isolating the sticky and nonsticky volume-fraction profiles from the linear combinations given in Eq. (17) and Eq. (18), we find the gas-phase sticky flux into the droplet,
$J_{\mathrm{s}\mathrm{g}}(R)$ $=$ $-\frac{D_{\mathrm{g}}}{R}[\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)]$ (20)
$-\frac{D_{\mathrm{g}}}{\ell}\frac{k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)-k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}.$
An identical expression for $J_{\mathrm{ns}\mathrm{g}}(R)$ is obtained upon replacing $\mathrm{s}\leftrightarrow\mathrm{ns}$. For growing droplets, the ratio $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}/\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ between sticky and nonsticky volume fractions in the gas phase just outside the droplet, is smaller than the one dictated by the reactions, $k_{\mathrm{ns}\to\mathrm{s}}/k_{\mathrm{s}\to\mathrm{ns}}=\varphi_{\mathrm{s}}^{\infty}/\varphi_{\mathrm{ns}}^{\infty}$. Thus, the second term in Eq. (20), resulting from the reactions, joins the first term arising from passive Ostwald ripening (Eq. (8)) in bringing material into the droplet, and vice versa for shrinking droplets. This fact is illustrated in Fig. 2, and it is the possible source of accelerated Ostwald ripening.
Figure 2: Illustrative concentration profiles between a small and a large droplet in the passive and active cases, where $\Delta\varphi_{\mathrm{s}}=\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)$ is the excess saturation of sticky particles in the dilute phase. In the passive case, close to each of the droplets, the concentration profile is hyperbolic (dashed black curve), bringing material to the large droplet and away from the small one, depending on the dilute phase concentration between the droplets. In the active case, reactions create nonsticky material, thus overall lowering the sticky volume fraction in the dilute phase. However, the reactions may make the sticky profile steeper in the vicinity of droplets, thus increasing the flux of material from the small to the large droplet. In both cases, the indicated proportionalities for $\Delta\varphi_{\mathrm{s}}$ hold for both the smaller and larger droplet, up to a change of sign for the large droplet. The full expressions for the volume fraction profiles can be found in the SI Appendix, Sec. II.
As mentioned above, linearizing Eq. (13) in the dense phase is much more intricate, and we consign these calculations to the SI Appendix, Sec. II. In contrast to the dilute phase, the dense phase is almost entirely composed of sticky proteins. This means that reactions inside the droplet hinder its growth, as their net effect is to convert sticky material into the nonsticky form, $s_{\mathrm{l}}\equiv s(\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}})>0$. (We first mention that a phenomenon akin to Turing instability [40] might occur for certain diffusive coupling values between sticky and nonsticky species. In this case, droplets may not be stable once $R(t)$ is larger than the characteristic length for the patterns. We do not explore this phenomenon further, and we assume that the parameter values are such that it is avoided, and hence the volume fraction profiles inside the droplets are stable.) For small droplets, the amount of converted material per time is proportional to droplet volume. However, once the droplet is larger than the distance a particle typically diffuses between reactions inside the droplet, the reactions equilibrate at the core of the droplet, [^4] We assume that the equilibrated volume fraction of sticky species, given by $s(\phi_{\mathrm{s}\mathrm{l}}(0;R),\phi_{\mathrm{ns}\mathrm{l}}(0;R))=0$ and $\phi_{\mathrm{s}\mathrm{l}}(0;R)+\phi_{\mathrm{ns}\mathrm{l}}(0;R)=\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}(R)+\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}(R)$ is still in between the binodal and spinodal lines of the free energy, so bubbles do not form [9]; see SI Appendix, Sec. II. so the net loss of sticky material occurs only at its interface and hence the total conversion rate is proportional to droplet surface area. By contrast, Ostwald ripening, which opposes this loss of material, occurs with a rate proportional to droplet radius, whereas the gas-phase reactions contribute additional ripening at a rate proportional to the area as well. These competitions produce a set of nonequilibrium behaviors which are explored below and summarized in Fig. 3.
We conclude this section by calculating the droplet growth rate. By equating Eq. (5) and Eq. (12), we find to leading order in small gas-phase volume fractions,
$\frac{k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)-k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}-\frac{s_{\mathrm{l}}g(R/\ell)}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}$
$=\frac{\ell}{R}[\varphi^{\infty}_{\mathrm{ns}}-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)]$ , (21)
from which we can infer $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ [37]. Details of this calculations, as well as the functional form of $g(x)$ (originating from the dense-phase calculations), can be found in SI Appendix, Sec. II. On the other hand, expressing Eq. (5) and Eq. (12) for the total volume fraction field $\phi_{\mathrm{s}}+\phi_{\mathrm{ns}}$, we find
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R}\left[\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)\right].$ (22)
Eq. (22) is a key result — upon comparison with Eq. (9), it emphasizes that regardless of the activity, this system is bound by the scaling laws of conventional Ostwald ripening. As mentioned previously, this is guaranteed by the conservation law of $\phi_{\mathrm{s}}+\phi_{\mathrm{ns}}$, a volume fraction field which is not directly affected by activity. Therefore, in our search for acceleration, we understand that the scaling laws expected from Ostwald ripening are the best case scenario, requiring small $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, and our only hope is to increase the rate coefficient. Otherwise, in the “worst case scenario” of large $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, Ostwald ripening might be completely prevented. The alternative possible outcomes we capture below include size control [3] and arrested ripening (that is, no shrinkage of small droplets) [6].
With this we have laid out our theory, and are prepared to list our results. In Sec. II.4, only the lengthscale $\lambda$ associated with surface tension was relevant, and was assumed to be much smaller than the typical droplet size $R$ during ripening. Here, $\ell$ of Eq. (19) is another, larger lengthscale to which $R$ must be compared. Thus, we split our discussion into small ($R\ll\ell$) and large ($R\gg\ell$) droplets. Detailed calculations can be found in SI Appendix, Sec. II.
IV Results
With the theoretical setup of Sec. III, we proceed to obtain explicitly the dynamics of a focal droplet. The reactions complicate the standard Ostwald ripening calculations of Sec. II [24]. First, the distance a particle diffuses before reacting splits droplet growth into two regimes, which we study in separate sections. Furthermore, the reactions within the liquid hinder Ostwald ripening, requiring special attention to identify the parameter values for which growth is permitted. Recent reviews on the known possible outcomes can be found in Refs. [41, 1].
IV.1 Small droplet
For small droplets, we compute the undetermined $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ using Eq. (21) to leading order in small dilute-phase density $\varphi^{\infty}\ll 1$ and radius $R/\ell\ll 1$, and keep the leading order term containing the net reaction rate inside the droplet, $s_{\mathrm{l}}$,
$\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)$ $=$ $\varphi^{\infty}_{\mathrm{ns}}-\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}[\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)]\frac{R}{\ell}$ (23)
$+\frac{1}{3}\frac{s_{\mathrm{l}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}\left(\frac{R}{\ell}\right)^{2},$
where $g(x)=x/3$ in this limit. Equipped with $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)$, Eq. (22) becomes
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R}\left(\Delta\varphi_{\mathrm{s}}-\frac{\lambda}{R}\right)\left(1+\frac{R}{\ell}\right)-\frac{s_{\mathrm{l}}R}{3}.$ (24)
where we defined the excess saturation $\Delta\varphi_{\mathrm{s}}=\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}$ and $\lambda$ is the Laplace pressure correction of the sticky volume fraction in the gas; see Sec. II.1. The growth-enhancing factor $1+R/\ell$ arises from the reactions in the dilute phase, whereas the last term, hindering growth, originates from the reactions in the dense phase. Indeed small droplets, $R\ll\lambda/\Delta\varphi_{\mathrm{s}}$, vanish. In the absence of reactions ($\ell\to\infty$, $s_{\mathrm{l}}=0$), Eq. (24) exactly coincides with the passive Lifshitz-Slyozov growth law [24]. We now elaborate on the results for the active case, which are summarized in Fig. 3.
If $s_{\mathrm{l}}$ is sufficiently large, indefinite growth and ripening are prevented. This follows because sticky proteins are lost at a rate proportional to the volume of the droplet, whereas they join from the dilute phase at a rate proportional to the radius. Thus, even for small $R$, large $s_{\mathrm{l}}$ may be sufficient to counteract the ripening. In this case, the radius $R^{*}=(3D\Delta\varphi_{\mathrm{s}}/s_{\mathrm{l}})^{1/2}$ is a stable point of Eq. (24), where we kept leading order in $\lambda/R^{*}\ll 1$ [3]. This implies that smaller droplets will grow and larger ones will shrink until $R^{*}$ is reached. [^5] Note that as droplets grow, the excess saturation $\varphi_{\mathrm{s}}^{\infty}$ decreases (cf. Eq. (10)), lowering further the terminal radius $R^{*}$. Therefore, $R^{*}=(3D\Delta\varphi_{\mathrm{s}}/s_{\mathrm{l}})^{1/2}$ is in fact a transcendental equation for $R^{*}$. This size control has been reported and discussed in depth, e.g., in Ref. [3] and observed in biological systems [4, 5]. In this case, the large droplet regime is inaccessible.
On the other hand, if the above terminal radius is too large, $R^{*}\gg\ell$, then size control does not occur before the droplets reach the $R\gg\ell$ regime we discuss below. This would be the case if the reactions occur too slowly, $s_{\mathrm{l}}<D_{\mathrm{g}}\Delta\varphi_{\mathrm{s}}/\ell^{2}$ (to leading order in $\lambda/\ell\ll 1$). In that case, the second term in Eq. (24) can be neglected. We comment that the additional factor $1+R/\ell$, originating from the reactions in the dilute phase, accelerates the growth of droplets of size $R\lesssim\ell$, though to a subleading extent. Therefore, in total, we recover standard Ostwald ripening (without meaningful acceleration) which continues until the growth dynamics transitions to the large droplet regime, $R\gg\ell$, discussed next.
IV.2 Large droplet
Figure 3: Phase diagram of droplet dynamics as a function of total dense phase reaction rate and droplet size. Reactions in the dilute phase may accelerate Ostwald ripening (cf. Fig. 2), whereas reactions in the dense phase hinder droplet growth and ripening by converting particles from sticky to nonsticky. Small droplets with $R\ll\lambda/\Delta\varphi_{\mathrm{s}}$ always tend to shrink due to large Laplace pressure. Other regions are approximately bounded by the inter-reaction diffusion length $\ell$. For very fast dense phase reactions, $s_{\mathrm{l}}\gg D_{\mathrm{g}}\Delta\varphi_{\mathrm{s}}/\ell^{2}$, size control is inevitable. Otherwise, droplets can reach the regime $R\gg\ell$, where: For intermediate dense phase reaction rate, $(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\lambda/\ell\ll s_{\mathrm{l}}\ll D\Delta\varphi_{\mathrm{s}}/\ell^{2}$, a negative effective surface tension prevents small droplets from shrinking. For slow dense phase reaction rate, $s_{\mathrm{l}}\ll(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\lambda/\ell$, Ostwald ripening occurs with an accelerated rate. Parameters: $R$, droplet radius; $s_{\mathrm{l}}$, total dense phase reaction rate; $\Delta\varphi_{\mathrm{s}}$, dilute phase volume fraction relative to the coexistence value; $\lambda$, characteristic width of a droplet interface; $\ell$, characteristic distance between reactions; $k_{i\to j}$, dilute-phase switching rates, $D_{\mathrm{g}}$, dilute-phase diffusion constant.
For larger droplets, solving Eq. (21) for $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ in the limit $R/\ell\gg 1$ yields,
$\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)$ $=$ $\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)+\frac{s_{\mathrm{l}}}{k_{\mathrm{ns}\to\mathrm{s}}}$
$+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\left[\left(1+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\right)\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}-\frac{2s_{\mathrm{l}}}{k_{\mathrm{s}\to\mathrm{ns}}}\right]\frac{\ell}{R},$
and inserting the result above in Eq. (22) results in
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{D^{\mathrm{eff}}}{R}\left(\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}-\frac{\lambda^{\mathrm{eff}}}{R}\right),$ (26)
where [^6] A reactive lengthscale similar to Eq. (19) emerges for the dense phase. To obtain Eqs. (28) and (29), we assumed that it is the same lengthscale $\ell$ in both phases. If these lengths are different, our results do not change qualitatively; rather, $s_{\mathrm{l}}$ will be multiplied by a finite factor in each expression. We address this detail in SI Appendix, Sec. II.
$D^{\mathrm{eff}}$ $=$ $D_{\mathrm{g}}\left(1+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\right),$ (27)
$\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$ $=$ $\Delta\varphi_{\mathrm{s}}-\frac{s_{\mathrm{l}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}},$ (28)
$\lambda^{\mathrm{eff}}$ $=$ $\lambda-\frac{2s_{\mathrm{l}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}\ell+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}\ell.$ (29)
Notably, the dilute-phase reactions have successfully increased the effective diffusion coefficient by a factor $1+k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}}$. The effective supersaturation (Eq. (28)), however, is smaller than the actual residual saturation $\Delta\varphi_{\mathrm{s}}$ due to net conversion of sticky material to nonsticky inside the droplet. This implies that, if Ostwald ripening occurs, it will stop earlier than in the passive case due to material exhaustion (Eq. (10)).
We proceed to discuss the possible growth outcomes implied by Eq. (26), summarized in Fig. 3, by inspecting the effective Laplace pressure correction coefficient, $\lambda^{\mathrm{eff}}$. Understanding it is intricate: Growth driven by Laplace pressure, as mentioned above, can only occur with a rate proportional to the radius, while reactions on either side of the interface (consuming and creating sticky protein inside and outside the droplet, respectively) occur with rates proportional to the area. Since now $R\gg\ell$, the growth matching condition, Eq. (21), requires the reactions to counteract one another, thus canceling each other. Hence, the effects described below arise from the next-order corrections; see SI Appendix, Sec. II.
When the reactions inside the droplet are sufficiently fast, $s_{\mathrm{l}}\gg(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\lambda/\ell$, we may neglect the contribution from the Laplace pressure, the first term in Eq. (29). Thus, temporarily ignoring the third term in Eq. (29), the reactions cause an effective negative Laplace pressure correction coefficient, $\lambda^{\mathrm{eff}}=-2\ell s_{\mathrm{l}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})$. As a result, the droplets do not undergo ripening per se, i.e., smaller droplets do not shrink. Instead, all droplets that pass the $R\gg\ell$ threshold are expected to grow until the extensive, stable radius $R^{\dagger}=\lambda^{\mathrm{eff}}/\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}\sim V^{1/3}$ is reached. For this stable radius, the effective supersaturation becomes negative. Thus, now including the third term of Eq. (29), we see that it will further contribute to the already negative $\lambda^{\mathrm{eff}}=-\ell[s_{\mathrm{l}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})+(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})|\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}|]$. As growth continues, material is being taken up from the dilute phase, so $\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$ becomes even more negative so $|\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}|$ increases. This ripening arrest has been recently reported in Ref. [6]; we will not investigate it here further.
We now discuss the sought-after regime where Ostwald ripening is accelerated. For slow reactions inside the droplet, $s_{\mathrm{l}}\ll(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\lambda/\ell$, the Laplace pressure dominates as the second term of Eq. (29) is negligible. Thus, the Laplace pressure correction coefficient is positive, and ripening occurs as expected. Here, $\lambda^{\mathrm{eff}}=\lambda+\ell(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$, where the effective supersaturation, $\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$, as in standard Ostwald ripening, is positive. Since $\ell/\lambda$ is a large number, the contribution from the small $\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$ may be transiently important. During a transient where $\lambda\ll\ell(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$, we find $\mathrm{d}R/\mathrm{d}t=(D^{\mathrm{eff}}\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}/R)[1+(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})(\ell/R)]\simeq D^{\mathrm{eff}}\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}/R$, meaning that $R^{2}\sim t$ until the effective supersaturation lowers sufficiently due to ripening.
With the progression of ripening, once $\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$ decreases sufficiently, we reproduce the standard $\lambda^{\mathrm{eff}}=\lambda$, and hence also the scaling of Lishitz-Slyozov [24], but with a larger diffusion constant compared to the passive Eq. (11),
$\langle R(t)\rangle^{3}-\langle R(0)\rangle^{3}_{\mathrm{a}}=\frac{4}{9}D^{\mathrm{eff}}\lambda t.$ (30)
Note that due to the transient for $R\ll\ell$ and possibly altered nucleation rates [42, 1], the intercept $\langle R(0)\rangle^{3}_{\mathrm{a}}$ might be different from the one in the passive case. Thus, the average droplet volume grows by a factor $D^{\mathrm{eff}}/D_{\mathrm{g}}=1+k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}}$ faster compared to the passive case. Interestingly, the faster the switching from sticky to nonsticky compared to the reverse process, the faster is the ripening rate. This increase of $D^{\mathrm{eff}}$ occurs at the expense of having a lower sticky supersaturation, Eq. (16). The fact that this is the only regime and mechanism by which Ostwald ripening can be accelerated is the main result of the paper.
IV.3 Realizability of accelerated ripening
With our theoretical results for accelerated ripening laid out, we assess their experimental relevance. Pragmatically, it is not trivial to realize the above constellation of assumptions. For clarity, we recap our approximations and explain the sacrifices made so as to achieve acceleration. Despite these shortcomings, we will argue for the feasibility to accelerate ripening and present the experimental results of Ref. [15] as evidence for a workable protocol.
The first pitfall is size control which may occur in the small droplet regime, $R\ll\ell$. Suppose the reactions inside the droplet occur unhindered as in the dilute phase, meaning that, to leading order, $s_{\mathrm{l}}\simeq k_{\mathrm{s}\to\mathrm{ns}}$ (where we have taken $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}\simeq 1$ and $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\simeq 0$). In that case, $R^{*}\sim(D_{\mathrm{g}}\Delta\varphi_{\mathrm{s}}/k_{\mathrm{s}\to\mathrm{ns}})^{1/2}\sim\ell\Delta\varphi_{\mathrm{s}}^{1/2}$, meaning that $R^{*}/\ell\ll 1$ and size control is bound to occur and thus the $R\gg\ell$ regime is inaccessible. To avoid this constraint, the reactions in the dense phase must occur with a reduced rate by excluding the enzymes or a fuel for reactions from the droplet. This can be achieved if, e.g., these catalyzers are bulky molecules and the droplets are very dense, or the sticky protein and the catalyzers have opposite hydrophilicities [15], or more generally if the sticky proteins inside the droplet apply a large entropic pressure against intruders [43]. We thus suppose that one of these mechanism causes the total reaction rate inside the droplet to be $s_{\mathrm{l}}<k_{\mathrm{s}\to\mathrm{ns}}\Delta\varphi_{\mathrm{s}}$, so $R$ may become comparable to or larger than $\ell$.
The second trouble may be the finite size of biological systems. For reasonable diffusivities $D_{\mathrm{g}}\sim 10^{-1}-10^{2}\mathrm{\mu m}^{2}\mathrm{sec}^{-1}$ and conversion rates $k_{i\to j}\sim 10^{-3}-10^{-1}\mathrm{sec}^{-1}$ [19], the distance between reactions may therefore vary between $\ell\sim 10^{-1}-10^{2}\mathrm{\mu m}$. The upper limit is larger than the typical eukaryotic cell size ($10\mathrm{\mu m}$), meaning that the accelerated ripening regime may not be attainable for slow reactions in biological systems. Therefore, we assume that the system size is sufficient to attain the $R\gtrsim\ell$ regime.
Once the regime $R>\ell$ is reached, we note that growth without ripening can be somewhat easily avoided. Since size control is overcome, $s_{\mathrm{l}}\sim D_{\mathrm{g}}\Delta\varphi_{\mathrm{s}}/\ell^{2}\sim(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\Delta\varphi_{\mathrm{s}}$, where $\Delta\varphi_{\mathrm{s}}\ll 1$. It is thus also possible to further require $s_{\mathrm{l}}\lesssim(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\lambda/\ell$ for $\lambda/\ell\ll 1$. Under these conditions, accelerated ripening is achieved.
However, the third problem is how long does accelerated ripening continue? For our approach to hold in the absence of reactions (left panel of Fig. 1), we assumed small $\Delta\varphi_{\mathrm{s}}=\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}$. To achieve meaningful acceleration, we need the deactivation reaction to be fast, $k_{\mathrm{s}\to\mathrm{ns}}\gg k_{\mathrm{ns}\to\mathrm{s}}$. This means, that in the active case, where $\Delta\varphi_{\mathrm{s}}=k_{\mathrm{ns}\to\mathrm{s}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})-\phi_{\mathrm{s}\mathrm{g}}^{\circ}$, supersaturation may not be large enough to produce an extended period of acceleration until ripening stops (let alone producing $\Delta\varphi_{\mathrm{s}}>0$ to allow ripening to occur in the first place). Moreover, the relevant supersaturation for the $R\gg\ell$ regime is in fact the even smaller $\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$ of Eq. (28) (where $s_{\mathrm{l}}\sim k_{\mathrm{s}\to\mathrm{ns}}\Delta\varphi_{\mathrm{s}}$ with the above arguments). Therefore, not only must $s_{\mathrm{l}}$ be small, but $\phi_{\mathrm{s}\mathrm{g}}^{\circ}$ must be very small, too, so the solution remains effectively supersaturated.
In total, accelerated ripening can be realized in scenarios where there are almost no reactions inside the dense phase (so growth will not be hindered), the gas-phase reactions are not too slow (so they occur on lengthscales smaller than the system size), and the sticky species is almost insoluble (so the active solution would indeed be supersaturated).
Figure 4: Active acceleration of Ostwald ripening. The solid red curve is the theoretical prediction of the acceleration factor, $1+k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}}$. The experimental data points are obtained from Ref. [15] in a case where the fuel concentration was kept constant up to an error. No fitting parameters were employed.
An encouraging artificial chemical system that was shown to exhibit accelerated ripening according to this mechanism is reported in Ref. [15]. There, the sticky species is a strongly hydrophobic molecule which, upon exposure to water, breaks down to form a nonsticky (hydrophilic) molecule. Introducing a specialized fuel molecule allows conversion of the hydrophilic species through a sequence of reactions into the hydrophobic one. Since the sticky species creates a hydrophobic condensate impenetrable to water, its degradation does not occur at all inside droplet ($s_{\mathrm{l}}=0$). At the same time, since the degradation rate, $k_{\mathrm{s}\to\mathrm{ns}}$, is faster than the formation in the dilute phase, $k_{\mathrm{ns}\to\mathrm{s}}$, indeed an acceleration factor of up to a $100$-fold is reported [15].
In Fig. 4 we plot our theoretical prediction along with the experimental results of Ref. [15], [^7] These three datapoints are obtained from experiments in which the fuel concentration was kept constant. Seven more datapoints are available in Ref. [15] from experiments where all of the fuel was introduced at the beginning, so during ripening only degradation of the hydrophobic species takes place. In that case, the accelerated ripening is enabled by accelerated shrinkage of small droplets. Our theory is quasistatic, and hence does not apply to this regime. with no fitting parameters. We estimate the reactions rates from measurements and numerical estimates available in the supplemental information of Ref. [15]. The acceleration factor is the ratio between slopes of the growth law of $R^{3}$ in the active and passive cases. Since the hydrophobic species does not exist without the fuel molecule, the growth rates for the passive cases are estimated via the Lifshitz-Slyozov law using measurements of surface tensions and solubilities. The agreement is not perfect, likely due to the approximations involved in estimating the passive growth rate, as well as $k_{\mathrm{ns}\to\mathrm{s}}$, and the protocol of periodic addition of the fuel that drives $k_{\mathrm{ns}\to\mathrm{s}}$. At the same time, the overall acceleration trend is reassuringly reproduced. Ref. [15] thus provides evidence that this singular means of accelerating ripening is indeed experimentally realizable.
V Discussion
In this work, we theoretically studied the possibility of accelerating Ostwald ripening via chemical activity. We employed standard tools from linear-irreversible thermodynamics, Lishitz-Slyozov theory, and their extensions to include reactions to obtain explicit expressions for the diffusion-limited growth of droplets under ripening. We found that reactions in the sticky-particle-poor phase increase the flux of material out of small droplets and into large droplets. Indeed, if growth-hindering sticky-to-nonsticky reactions inside droplets can be avoided, we find acceleration of Ostwald ripening by an arbitrarily large factor, but with the same $R\sim t^{1/3}$ scaling imposed by particle conservation. Owing to the generality of our theory, we learn by elimination that this scenario in which reactions only occur in the dilute phase is indeed the only means to achieve accelerated Ostwald ripening. Our theoretical claims are qualitatively supported by recent experiments in an artificial chemical system [15]. Therefore, acceleration of ripening is another feat of active processes, joining the current ranks of size control [3, 4, 2, 5, 1], arrested ripening [6], Turing instabilities [7, 8], and bubbly phases [9].
The generality of our theory has been made possible through a series of simplifying assumption. Notably, we assumed that droplets are sparse, slowly growing, and that the source of activity does not modify surface thermodynamics. While these approximations are often reasonable, to actually achieve meaningful accelerated ripening is not trivial, requiring three components: (1) Reactions must be prevented inside the droplets. (2) The interconversion reactions must preferably form nonsticky species ($k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}}>1$). (3) The overall concentration of material must be high enough that at chemical equilibrium in the dilute phase, the sticky species is supersaturated; otherwise droplets will not form. The experimental system of Ref. [15] follows all these assumptions to the letter, although the fact that, to our knowledge, it is the only instance where an acceleration by these means was obtained points to the challenge of realizing these conditions.
Our theory has ignored droplet diffusion and coalescence — another means with which droplet growth can occur. Biomolecular condensates typically do not have a large surface tension, so growth by coalescence might in fact be the dominant effect driving the growth of natural biocondensates in cells [18]. To complicate matters further, growth by coalescense also exhibits $R\sim t^{1/3}$ scaling [25, 26], so distinguishing among growth by ripening from coalescence typically requires dynamic microscopy. Clearly, the interplay of reactions with droplet coalescence is a relevant future direction, as is the competition between the two growth mechanisms in active systems. We also ignored many other cellular complexities which interact with condensates [1]. In the context of the pyrenoid, these include membrane wetting mediated by molecular tethers [44] and the fact that two interacting components are required for droplet formation (Rubisco and EPYC1), rather than just a single “sticky” particle [23].
On a theoretical level, our work is a manifestation of an overall theme concerning the ability to accelerate thermodynamic relaxation processes. One extensively studied example is the Mpemba effect, where an out-of-equilibrium system’s initial condition is chosen so it will undergo relaxation along a faster trajectory [45]. Other accelerated phenomena include shortcuts to adiabaticity [46, 47, 48], escape by active particles [49, 50], and search processes under stochastic resetting [51, 52, 53]. In this context, it is important to stress that not all types or degrees of acceleration are physically achievable. Activity typically breaks time-reversal symmetry, as in our case, but any accelerated process is still constrained by remaining unbroken conservation laws.
On a practical level, to our knowledge, the acceleration method we report has only been realized in a chemical system [15]. Nevertheless, we posit that realizing accelerated Ostwald ripening in biocondensates is plausible. For example, advances in synthetic biology in principle allow the design of a phosphorylation-dependent switchable protein (like EPYC1) and a cognate kinase and phosphatase that are excluded from droplets [54], thus preventing conversion reactions in the dense phase. If rapid coarsening can be turned on or off at will, e.g., by optogenetics [55], the resulting size control of condensates would enable novel approaches to metabolic control [16, 17].
To conclude, it is well established that living systems consume energy to carry out biological processes [56]. By driving molecular machinery “uphill” against chemical potential differences and thus breaking detailed balance, these active processes oppose relaxation to thermodynamic equilibrium. On the other hand, when relaxation to equilibrium is desired, it may be beneficial to accelerate the process. This fact is showcased by the ubiquity of enzymatic catalysis — a ubiquitous, non-energy-consuming means to accelerate biological reactions [57]. Accelerating processes such as phase separation as studied here but also membrane fission [58, 59] and fusion [60, 61, 62], DNA hybridization [63], and ligand-receptor binding [64] may be advantageous as well. It remains an exciting yet understudied question when and where cells use energy to accelerate thermodynamic relaxation.
Acknowledgements
We thank Job Boekhoven and Alejandro Martínez-Calvo for illuminating discussions. B.S. acknowledges support from the Princeton Center for Theoretical Science. This project has been made possible in part by grant number DAF2024-342781 from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation, and NIH Grant R01GM140032.
Supplementary Material: Accelerated Ostwald ripening by chemical activity
SI Thermodynamic prerequisites and the Flory-Huggins free energy
Most of the generic conclusions drawn in the main text for the equilibrium properties of phase-separating systems are motivated by the Flory-Huggins (FH) model [65, 66]. In this section, we first present the general thermodynamic principles underlying phase separation of a binary fluid (sticky and inert solvent). We proceed to connect the surface tension to the mesoscopic field theory considered in the main text. We then show concrete results for the FH free energy to provide intuition and to further support the assumptions concerning the volume fractions of each phase. We conclude by extending all these results to the case of a ternary fluid (sticky, nonsticky, and inert solvent) which is the case studied in the main text.
SI.1 Thermodynamics of coexistence
First, we outline the general thermodynamic principles governing phase-separating systems. We consider a material possessing a free energy $F$, which is a function of temperature $T$, volume $V$, and number of particles $\{N_{i}\}$, where $i=\mathrm{s}$ denotes the sticky particles and $i=\mathrm{o}$ the remaining inert and indistinguishable biomolecules of the cell. (We will add in the nonsticky species, $i=\mathrm{ns}$, in Sec. SI.4 below). We suppose that the total local concentration of biomolecules (mainly proteins and RNAs) in the cytoplasm is fixed by strong osmotic constraints, hence, $\varrho=(N_{\mathrm{s}}+N_{\mathrm{o}})/V=\mathrm{const}$. This imposes an effective incompressibility condition on the dissolved molecules,
$\phi_{\mathrm{s}}+\phi_{\mathrm{o}}=1,$ (S1)
where $\phi_{i}=N_{i}/(V\varrho)$ is the volume fraction of the $i$th species. Thus, $\phi_{\mathrm{o}}$ is uniquely determined by $\phi_{\mathrm{s}}$, as well as $N_{\mathrm{o}}$ by $N_{\mathrm{s}}$ and $V$. Furthermore, we assume that the temperature $T$ is also held constant. Upon phase separation, only $V$ and $N_{\mathrm{s}}$ vary, and so we will suppress the dependence of the free energy $F$ on the constants $T$ and $\varrho$.
With the above setup, Euler’s homogeneous functions principle allows us to write the free energy as [27]
$F(V,N_{\mathrm{s}})=V\varrho k_{\mathrm{B}}Tf\left(\frac{N_{\mathrm{s}}}{V\varrho}\right),$ (S2)
where $f(\phi_{\mathrm{s}})$ is the total normalized free-energy per molecule (including both sticky and inert species). The sticky chemical potential and pressure are given by
$\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}})}{k_{\mathrm{B}}T}$ $=$ $\frac{1}{k_{\mathrm{B}}T}\left(\frac{\partial F}{\partial N_{\mathrm{s}}}\right)_{V}=f^{\prime}(\phi_{\mathrm{s}}).$ (S3a)
$\frac{P(\phi_{\mathrm{s}})}{\varrho k_{\mathrm{B}}T}$ $=$ $-\frac{1}{\varrho k_{\mathrm{B}}T}\left(\frac{\partial F}{\partial V}\right)_{N_{\mathrm{s}}}=\phi_{\mathrm{s}}f^{\prime}(\phi_{\mathrm{s}})-f(\phi_{\mathrm{s}}),$ (S3b)
where $f^{\prime}\equiv\partial f/\partial\phi_{\mathrm{s}}$ for shorthand.
Phase separation is possible whenever the free energy has more than one stable volume fraction. We denote the coexistence volume fraction of the dense phase (sticky-rich, “liquid”) by $\phi_{\mathrm{s}\mathrm{l}}^{\circ}$ and of the dilute phase (sticky-poor, “gas”) by $\phi_{\mathrm{s}\mathrm{g}}^{\circ}$. When two phases in contact are in equilibrium, the sticky chemical potential and the pressure of the two phases equate [27],
$\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{l}}^{\circ})$ $=$ $\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ}),$ (S4a)
$P(\phi_{\mathrm{s}\mathrm{l}}^{\circ})$ $=$ $P(\phi_{\mathrm{s}\mathrm{g}}^{\circ}).$ (S4b)
These two equations determine the volume fractions of each phase. Furthermore, a known total amount of sticky material $N_{\mathrm{s}}$ allows one to determine the volume fraction of the dense phase $\zeta_{\mathrm{l}}=1-\zeta_{\mathrm{g}}$ when phase separation is complete, $\zeta_{\mathrm{l}}\phi_{\mathrm{s}\mathrm{l}}^{\circ}+(1-\zeta_{\mathrm{l}})\phi_{\mathrm{s}\mathrm{g}}^{\circ}=N_{\mathrm{s}}/(V\varrho)$.
In the main text, the liquid phase is assumed to be finite sized and spherical with radius $R$. As a result, we must account for the surface tension $\sigma$, which leads to a curvature-induced Laplace pressure [28], $2\sigma/R$. Therefore, The coexistence volume fractions modify to $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}$ and $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$, which satisfy
$\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}})$ $=$ $\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}),$ (S5a)
$P(\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}})$ $=$ $P(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}})+\frac{2\sigma}{R}.$ (S5b)
Assuming the droplet is large (so Laplace pressure is small), we approximate $\phi_{\mathrm{s}\alpha}^{\mathrm{L}}\simeq\phi_{\mathrm{s}\alpha}^{\circ}+\lambda_{\mathrm{s}\alpha}/R$ in each phase $\alpha=\mathrm{l},\mathrm{g}$ (where $\{\lambda_{\mathrm{s}\alpha}\}$ are Laplace-pressure correction constants), and expand Eqs. (S5) for small $\lambda_{\mathrm{s}\alpha}/R$. This yields Eqs. (S4) for the leading order, and the following next-order equations:
$f^{\prime\prime}(\phi_{\mathrm{s}\mathrm{l}}^{\circ})\lambda_{\mathrm{s}\mathrm{l}}$ $=$ $f^{\prime\prime}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})\lambda_{\mathrm{s}\mathrm{g}},$ (S6a)
$\phi_{\mathrm{s}\mathrm{l}}^{\circ}f^{\prime\prime}(\phi_{\mathrm{s}\mathrm{l}}^{\circ})\lambda_{\mathrm{s}\mathrm{l}}$ $=$ $\phi_{\mathrm{s}\mathrm{g}}^{\circ}f^{\prime\prime}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})\lambda_{\mathrm{s}\mathrm{g}}+\frac{2\sigma}{\varrho k_{\mathrm{B}}T}.$ (S6b)
The solution to these equations is
$\lambda_{\mathrm{s}\alpha}=\frac{1}{\phi_{\mathrm{s}\mathrm{l}}^{\circ}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}}\frac{2\sigma}{\varrho k_{\mathrm{B}}Tf^{\prime\prime}(\phi_{\mathrm{s}\alpha}^{\circ})}=\frac{2(\phi_{\mathrm{s}\alpha}^{\circ})^{2}\beta(\phi_{\mathrm{s}\alpha}^{\circ})\sigma}{\phi_{\mathrm{s}\mathrm{l}}^{\circ}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}}.$ (S7)
where, using Eq. (S3b), we identified the isothermal compressibility coefficient [36] $\beta(\phi_{\mathrm{s}})\equiv-(1/V)(\partial V/\partial P)_{N_{\mathrm{s}}}=1/[\varrho k_{\mathrm{B}}T\phi_{\mathrm{s}}^{2}f^{\prime\prime}(\phi_{\mathrm{s}})]$. Indeed, the higher the surface tension $\sigma$ and compressibility $\beta$ (the susceptibility of each phase to compression due to the extra Laplace pressure), the larger is the increase in sticky volume fraction due to Laplace pressure. In the derivations to follow, only the gas-phase sticky Laplace-pressure correction will appear explicitly, which is why we write for brevity $\lambda_{\mathrm{s}\mathrm{g}}\equiv\lambda$ in the main text.
To conclude, for completeness, we present the following common simplification of the above: In a typical system, the liquid phase is much denser than the gas phase, $\phi_{\mathrm{s}\mathrm{l}}^{\circ}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}\simeq\phi_{\mathrm{s}\mathrm{l}}^{\circ}$, the former is incompressible, $\beta(\phi_{\mathrm{s}\mathrm{l}}^{\circ})\simeq 0$, and the latter is ideal, $\beta(\phi_{\mathrm{s}\mathrm{g}}^{\circ})\simeq 1/(\varrho k_{\mathrm{B}}T\phi_{\mathrm{s}\mathrm{g}}^{\circ})$. Thus, the liquid typically has no Laplace-pressure correction, $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}\simeq\phi_{\mathrm{s}\mathrm{l}}^{\circ}$ whereas the gas follows the Ostwald-Freundlich equation [67],
$\frac{\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}}{\phi_{\mathrm{s}\mathrm{g}}^{\circ}}\simeq 1+\frac{1}{\varrho k_{\mathrm{B}}T\phi_{\mathrm{s}\mathrm{l}}^{\circ}}\frac{2\sigma}{R}\simeq\exp\left[\frac{2v_{\mathrm{m}}\sigma}{k_{\mathrm{B}}TR}\right],$ (S8)
where $v_{\mathrm{m}}=(\varrho\phi_{\mathrm{s}\mathrm{l}}^{\circ})^{-1}$ is the volume per sticky particle in the liquid phase.
SI.2 Surface tension
In our coarse-grained view of a phase-separating system, a given particle configuration is modeled via a volume-fraction field. To leading order in small-wavelength excitations, we promote the homogeneous free energy (Eq. (S2)) into a free energy functional,
$F[\phi_{\mathrm{s}}(\mathbf{r})]=\varrho k_{\mathrm{B}}T\int\mathrm{d}\mathbf{r}\left[f(\phi_{\mathrm{s}}(\mathbf{r}))+\frac{\kappa_{\mathrm{s}}}{2}|\boldsymbol{\nabla}\phi_{\mathrm{s}}(\mathbf{r})|^{2}\right],$ (S9)
where we have included the leading-order energetic penalty in volume-fraction gradients, whose magnitude is controlled by $\kappa_{\mathrm{s}}$. Equation (S9) coincides with Eq. (S2) for a homogeneous medium. Incompressibility (Eq. (S1)) is assumed at every position, $\phi_{\mathrm{s}}(\mathbf{r})+\phi_{\mathrm{o}}(\mathbf{r})=1$. In this section, we show how the macroscopic surface tension $\sigma$ can be related to the microscopic penalty $\kappa_{\mathrm{s}}$ [29, 68].
We compute the surface tension due to a flat interface between two semi-infinite bulks. (Any curvature might, in principle, result in an additional $\mathcal{O}(R^{-1})$ contribution to the surface tension, which is already subleading when inserted into the Laplace-pressure corrections to the volume fractions.) We denote the axis perpendicular to the interface as $x$, with the interface lying on $x=0$, and we arbitrarily designate the left side as the gas, $\phi_{\mathrm{s}}(-\infty)=\phi_{\mathrm{s}\mathrm{g}}^{\circ}$, and right side to be the liquid $\phi_{\mathrm{s}}(\infty)=\phi_{\mathrm{s}\mathrm{l}}^{\circ}$.
In equilibrium, the local chemical potential, $\delta F/\delta[\varrho\phi_{\mathrm{s}}(x)]$ (where $\varrho\phi_{\mathrm{s}}$ is the local particle density), is uniform everywhere so no particle transport would occur,
$f^{\prime}(\phi_{\mathrm{s}}(x))-\kappa_{\mathrm{s}}\frac{\mathrm{d}^{2}\phi_{\mathrm{s}}(x)}{\mathrm{d}x^{2}}=\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T},$ (S10)
where $\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})=\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{l}}^{\circ})$ is the chemical potential deep in each bulk. We multiply each side of the equation by $\mathrm{d}\phi_{\mathrm{s}}(x)/\mathrm{d}x$, and integrate from $\infty$- ∞ to $x$,
$f(\phi_{\mathrm{s}}(x))-\frac{\kappa_{\mathrm{s}}}{2}\left[\frac{\mathrm{d}\phi_{\mathrm{s}}(x)}{\mathrm{d}x}\right]^{2}-f(\phi_{\mathrm{s}\mathrm{g}}^{\circ})=\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}[\phi_{\mathrm{s}}(x)-\phi_{\mathrm{s}\mathrm{g}}^{\circ}].$ (S11)
The surface tension is the accumulated excess pressure due to the interface [29, 68] — i.e., the difference between the local pressure, $\varrho\phi_{\mathrm{s}}(x)\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})-\varrho k_{\mathrm{B}}T\{f(\phi_{\mathrm{s}}(x))+(1/2)\kappa_{\mathrm{s}}[\mathrm{d}\phi_{\mathrm{s}}(x)/\mathrm{d}x]^{2}\}$ (where $\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})$ is the uniform chemical potential, $\varrho\phi_{\mathrm{s}}(x)$ is the local density, and the second term is the local free energy density, with the three inserted in Eq. (S3b)), relative to the pressure without the interface, $\varrho\phi_{\mathrm{s}\mathrm{g}}^{\circ}\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})-\varrho k_{\mathrm{B}}Tf(\phi_{\mathrm{s}\mathrm{g}}^{\circ})=\varrho\phi_{\mathrm{s}\mathrm{l}}^{\circ}\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{l}}^{\circ})-\varrho k_{\mathrm{B}}Tf(\phi_{\mathrm{s}\mathrm{l}}^{\circ})$; cf. Eq. (S3b). Explicitly,
$\sigma=-\varrho k_{\mathrm{B}}T\int_{-\infty}^{\infty}\mathrm{d}x\Biggl{(}\left[\phi_{\mathrm{s}}(x)\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}-\left\{f(\phi_{\mathrm{s}}(x))+\frac{\kappa_{\mathrm{s}}}{2}\left[\frac{\mathrm{d}\phi_{\mathrm{s}}(x)}{\mathrm{d}x}\right]^{2}\right\}\right]-\left[\phi_{\mathrm{s}\mathrm{g}}^{\circ}\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}-f(\phi_{\mathrm{s}\mathrm{g}}^{\circ})\right]\Biggr{)}.$ (S12)
Expressing $(\kappa_{\mathrm{s}}/2)[\mathrm{d}\phi_{\mathrm{s}}(x)/\mathrm{d}x]^{2}$ via Eq. (S11), we simplify Eq. (S12) as
$\frac{\sigma}{\varrho k_{\mathrm{B}}T}$ $=$ $\sqrt{2\kappa_{\mathrm{s}}}\int_{-\infty}^{\infty}\mathrm{d}x\frac{\mathrm{d}\phi_{\mathrm{s}}(x)}{\mathrm{d}x}\left\{[f(\phi_{\mathrm{s}}(x))-f(\phi_{\mathrm{s}\mathrm{g}}^{\circ})]-\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}[\phi_{\mathrm{s}}(x)-\phi_{\mathrm{s}\mathrm{g}}^{\circ}]\right\}^{1/2}$ (S13)
$=$ $\sqrt{2\kappa_{\mathrm{s}}}\int_{\phi_{\mathrm{s}\mathrm{g}}^{\circ}}^{\phi_{\mathrm{s}\mathrm{l}}^{\circ}}\mathrm{d}a\left\{[f(a)-f(\phi_{\mathrm{s}\mathrm{g}}^{\circ})]-\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}(a-\phi_{\mathrm{s}\mathrm{g}}^{\circ})\right\}^{1/2},$
where we changed variables of integration from $x\in(-\infty,\infty)$ to $a=\phi(x)\in(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{s}\mathrm{l}}^{\circ})$. Thus we find the scaling $\sigma\sim\sqrt{\kappa_{\mathrm{s}}}$, which could have also been found from scaling arguments. The integral brings forth the dependence on the underlying energetics of the bulks.
SI.3 Flory-Huggins free energy
As a motivating example to support our choices of volume fractions and the general intuition in the main text, we specialize to the FH free energy [65, 66], which is widely used in modeling phase separation of simple liquids. It has two contributions; one coming from the entropy of mixing, and the other from sticky-sticky interactions,
$f(\phi_{\mathrm{s}})=\phi_{\mathrm{s}}\ln\phi_{\mathrm{s}}+(1-\phi_{\mathrm{s}})\ln(1-\phi_{\mathrm{s}})+\chi\phi_{\mathrm{s}}(1-\phi_{\mathrm{s}}),$ (S14)
where $k_{\mathrm{B}}T(-\chi)$ accounts for the attractive ($\chi>0$) energy between sticky particles. We assume that the molecular volumes of the species [1], as well as their polymeric lengths [69], are equal. The chemical potential and pressure are given by
$\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}})}{k_{\mathrm{B}}T}$ $=$ $\ln\phi_{\mathrm{s}}-\ln(1-\phi_{\mathrm{s}})+\chi(1-2\phi_{\mathrm{s}}),$ (S15a)
$\frac{P(\phi_{\mathrm{s}})}{\varrho k_{\mathrm{B}}T}$ $=$ $-\ln(1-\phi_{\mathrm{s}})-\chi\phi_{\mathrm{s}}^{2}.$ (S15b)
We find the volume fractions of each phase in coexistence. Phase separation occurs in this model for $\chi>2$ [69]. To model strong phase separation, $1-\phi_{\mathrm{s}\mathrm{l}}^{\circ},\phi_{\mathrm{s}\mathrm{g}}^{\circ}\ll 1$, we consider $\chi\gg 1$. Upon solving Eqs. (S4) for the FH free energy under the ansatz $\phi_{\mathrm{s}\mathrm{l}}^{\circ}\simeq 1$ and $\phi_{\mathrm{s}\mathrm{g}}^{\circ}\simeq 0$, we find
$\phi_{\mathrm{s}\mathrm{l}}^{\circ}$ $=$ $1-e^{-\chi},$ (S16a)
$\phi_{\mathrm{s}\mathrm{g}}^{\circ}$ $=$ $e^{-\chi},$ (S16b)
where corrections are of order $\chi e^{-2\chi}$, which is thus consistent with $\chi\gg 1$. Inserting these volume fractions in Eq. (S7), we find
$\lambda_{\mathrm{s}\mathrm{l}}=\lambda_{\mathrm{s}\mathrm{g}}\equiv\lambda=\frac{2\sigma e^{-\chi}}{\varrho k_{\mathrm{B}}T},$ (S17)
with, again, corrections of order $\chi e^{-2\chi}$. Interestingly, although the FH liquid phase is much less compressible than the gas phase, $\varrho k_{\mathrm{B}}T\beta(1-e^{-\chi})=e^{-\chi}\ll\varrho k_{\mathrm{B}}T\beta(e^{-\chi})=e^{\chi}$, the Laplace-pressure corrections are equal.
Lastly, we compute the surface tension $\sigma$ for FH in the large-$\chi$ limit. First, since $\kappa_{\mathrm{s}}$ is an energy penalty due to gradients, we expect $\kappa_{\mathrm{s}}=\chi\xi^{2}$, where $\xi$ is the effective lengthscale over which the microscopic sticky-sticky short-range potential is acting (whereas $\chi$ controls that potential’s magnitude). Thus, using Eq. (S12), we find for $\chi\gg 1$
$\frac{\sigma}{\varrho k_{\mathrm{B}}T}=\frac{\pi}{4\sqrt{2}}\xi\left[\chi-4\ln\left(\frac{4}{e}\right)\right],$ (S18)
where higher-order corrections are $\chi^{2}e^{-2\chi}$. Overall,
$\lambda=\frac{\pi}{2\sqrt{2}}\xi e^{-\chi}\left[\chi-4\ln\left(\frac{4}{e}\right)\right].$ (S19)
We see, therefore, that the large droplet approximation is consistent when $R\gg\chi\xi$, which is reasonable as droplets should be much larger than molecular lengthscales. Moreover, the scaling $\lambda\sim\xi\chi e^{-\chi}$ indicates that the surface tension increases more slowly with $\chi$ than the compressibility of the liquid phase decays. Thus, perhaps counterintuitively, Ostwald ripening occurs slower for systems which energetically tend more strongly to phase separate.
SI.4 Ternary mixtures
We now extend all formulas presented above to the case where a third, nonsticky component is present. The simplest conceptual way to think of the (inert) nonsticky species is as if one has labeled a given fraction of the inert proteome [70]. Thus, the total available protein concentration is still fixed at $\varrho=(N_{\mathrm{s}}+N_{\mathrm{ns}}+N_{\mathrm{o}})/V$, and the incompressibility condition (Eq. (S1)) is modified to include the nonsticky volume fraction $\phi_{\mathrm{ns}}$,
$\phi_{\mathrm{s}}+\phi_{\mathrm{ns}}+\phi_{\mathrm{o}}=1.$ (S20)
The free energy (Eq. (S2)), therefore, is rewritten as
$F(V,N_{\mathrm{s}},N_{\mathrm{s}})=V\varrho k_{\mathrm{B}}Tf\left(\frac{N_{\mathrm{s}}}{V\varrho},\frac{N_{\mathrm{ns}}}{V\varrho}\right),$ (S21)
where $f(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})$ is the amended normalized free-energy per molecule. The two chemical potentials ($i=\mathrm{s},\mathrm{ns}$) and pressure are now given by
$\frac{\mu_{i}(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})}{k_{\mathrm{B}}T}$ $=$ $\frac{\partial f}{\partial\phi_{i}}$ (S22a)
$\frac{P(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})}{\varrho k_{\mathrm{B}}T}$ $=$ $\phi_{\mathrm{s}}\frac{\partial f}{\partial\phi_{\mathrm{s}}}+\phi_{\mathrm{ns}}\frac{\partial f}{\partial\phi_{\mathrm{ns}}}-f(\phi_{\mathrm{s}},\phi_{\mathrm{ns}}).$ (S22b)
When the system phase separates, per the Gibbs phase rule [36], only three thermodynamic constraints are imposed on the four coexistence volume fractions $\{\phi_{i\alpha}^{\mathrm{L}}=\phi_{i\alpha}^{\circ}+\lambda_{i\alpha}/R\}$ (where $i=\mathrm{s},\mathrm{ns}$ and $\alpha=\mathrm{l},\mathrm{g}$),
$\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}})$ $=$ $\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}),$ (S23a)
$\mu_{\mathrm{ns}}(\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}})$ $=$ $\mu_{\mathrm{ns}}(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}),$ (S23b)
$P(\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}})$ $=$ $P(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})+\frac{2\sigma}{R}.$ (S23c)
All four volume fractions (as well as the four Laplace-pressure corrections) can be deduced by imposing material conservation across both phases when phase separation is complete, $\zeta_{\mathrm{l}}\phi_{\mathrm{s}\mathrm{l}}^{\circ}+(1-\zeta_{\mathrm{l}})\phi_{\mathrm{s}\mathrm{g}}^{\circ}=N_{\mathrm{s}}/(V\varrho)$ and $\zeta_{\mathrm{l}}\phi_{\mathrm{ns}\mathrm{l}}^{\circ}+(1-\zeta_{\mathrm{l}})\phi_{\mathrm{ns}\mathrm{g}}^{\circ}=N_{\mathrm{ns}}/(V\varrho)$, where $\zeta_{\mathrm{l}}=1-\zeta_{\mathrm{g}}$ is the volume fraction of the dense phase (total five equations with five unknowns $\{\phi_{i\alpha}^{\mathrm{L}}\}$ and $\zeta_{\mathrm{l}}$), or by flux matching during droplet growth (Eqs. (5) and (12) of the main text, for a total of five equations with five unknowns $\{\phi_{i\alpha}^{\mathrm{L}}\}$ and $\mathrm{d}R/\mathrm{d}t$).
We note that, since it is inert, the nonsticky species should not have a volume-fraction gradient penalty in the free energy functional. Thus, the free energy functional (Eq. (S9)) becomes
$F[\phi_{\mathrm{s}}(\mathbf{r}),\phi_{\mathrm{ns}}(\mathbf{r})]=\varrho k_{\mathrm{B}}T\int\mathrm{d}\mathbf{r}\left[f(\phi_{\mathrm{s}}(\mathbf{r}),\phi_{\mathrm{ns}}(\mathbf{r}))+\frac{\kappa_{\mathrm{s}}}{2}|\boldsymbol{\nabla}\phi_{\mathrm{s}}(\mathbf{r})|^{2}\right].$ (S24)
When estimating the surface tension at the interface between two semi-infinite bulks, the uniformity of both chemical potentials (cf. Eq. (S10)) must be required,
$\frac{\partial f}{\partial\phi_{\mathrm{s}}}\Biggr{|}_{\phi_{\mathrm{s}}(x),\phi_{\mathrm{ns}}(x)}-\kappa_{\mathrm{s}}\frac{\mathrm{d}^{2}\phi_{\mathrm{s}}(x)}{\mathrm{d}x^{2}}$ $=$ $\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T},$ (S25a)
$\frac{\partial f}{\partial\phi_{\mathrm{ns}}}\Biggr{|}_{\phi_{\mathrm{s}}(x),\phi_{\mathrm{ns}}(x)}$ $=$ $\frac{\mu_{\mathrm{ns}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}.$ (S25b)
Note that Eq. (S25b) implies that $\phi_{\mathrm{ns}}(x)$ simply spatially rearranges in accord with $\phi_{\mathrm{s}}(x)$, so the positional dependence of the former is strictly acquired from the latter, $\phi_{\mathrm{ns}}(x)\to\phi_{\mathrm{ns}}(\phi_{\mathrm{s}}(x))$. By multiplying each of Eqs. (S25) by $\mathrm{d}\phi_{i}(x)/\mathrm{d}x$, summing the two equations, and integrating the result from $\infty$- ∞ to $x$, we obtain
$f(\phi_{\mathrm{s}}(x),\phi_{\mathrm{ns}}(\phi_{\mathrm{s}}(x)))-\frac{\kappa_{\mathrm{s}}}{2}\left[\frac{\mathrm{d}\phi_{\mathrm{s}}(x)}{\mathrm{d}x}\right]^{2}-f(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})=\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}[\phi_{\mathrm{s}}(x)-\phi_{\mathrm{s}\mathrm{g}}^{\circ}]-\frac{\mu_{\mathrm{ns}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}[\phi_{\mathrm{ns}}(\phi_{\mathrm{s}}(x))-\phi_{\mathrm{ns}\mathrm{g}}^{\circ}],$ (S26)
Via a similar mechanism to Sec. SI.2, the surface tension can be found as
$\frac{\sigma}{\varrho k_{\mathrm{B}}T}=\sqrt{2\kappa_{\mathrm{s}}}\int_{\phi_{\mathrm{s}\mathrm{g}}^{\circ}}^{\phi_{\mathrm{s}\mathrm{l}}^{\circ}}\mathrm{d}a\Biggl{\{}[f(a,\phi_{\mathrm{ns}}(a))-f(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})]-\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}(a-\phi_{\mathrm{s}\mathrm{g}}^{\circ})-\frac{\mu_{\mathrm{ns}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}[\phi_{\mathrm{ns}}(a)-\phi_{\mathrm{ns}\mathrm{g}}^{\circ}]\Biggr{\}}^{1/2}.$ (S27)
To make the case with nonsticky particles concrete, we amend the FH free energy (Eq. (S14)) as
$f(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})=\phi_{\mathrm{s}}\ln\phi_{\mathrm{s}}+\phi_{\mathrm{ns}}\ln\phi_{\mathrm{ns}}+(1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}})\ln(1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}})+\chi\phi_{\mathrm{s}}(1-\phi_{\mathrm{s}}),$ (S28)
where the “labeling” of nonsticky particles results in an additional entropy contribution [70]. The internal energy term has remained unchanged as only the sticky particles are noninert. Once again, all species have been assumed to have the same molecular volumes [1] and number of monomers in the case of polymers [69]. The chemical potentials and pressure are given by
$\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})}{k_{\mathrm{B}}T}$ $=$ $\ln\phi_{\mathrm{s}}-\ln(1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}})+\chi(1-2\phi_{\mathrm{s}}),$ (S29a)
$\frac{\mu_{\mathrm{ns}}(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})}{k_{\mathrm{B}}T}$ $=$ $\ln\phi_{\mathrm{ns}}-\ln(1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}}),$ (S29b)
$\frac{P(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})}{\varrho k_{\mathrm{B}}T}$ $=$ $-\ln(1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}})-\chi\phi_{\mathrm{s}}^{2}.$ (S29c)
While only three thermodynamic constraints are given at coexistence (Eqs. (S23)), we can make progress in the large-$\chi$ regime for the FH-model nonetheless. First, via Eq. (S23b) we confirm that the nonsticky species remain equivalent to the inert solvent despite the labeling. Namely, the ratio between the nonsticky volume fraction ($\phi_{\mathrm{ns}}$) and the solvent one ($\phi_{\mathrm{o}}=1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}}$) remains fixed across both phases,
$\frac{\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}}{1-\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}-\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}}=\frac{\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}}{1-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}}.$ (S30)
With this assurance, we proceed to posit, once again, that $\phi_{\mathrm{s}\mathrm{l}}^{\circ}\simeq 1$ and $\phi_{\mathrm{s}\mathrm{g}}^{\circ}\simeq 0$ for $\chi\gg 1$, as the sticky protein is the one driving phase separation. Indeed, we once again find Eqs. (S16) and (S17), whereas the nonsticky volume fractions are found to satisfy $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}=\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(e^{-\chi}+\lambda/R)$, so
$\phi_{\mathrm{ns}\mathrm{l}}^{\circ}=\phi_{\mathrm{ns}\mathrm{g}}^{\circ}e^{-\chi}$ (S31)
and
$\lambda_{\mathrm{ns}\mathrm{l}}=\lambda_{\mathrm{ns}\mathrm{g}}e^{-\chi}+\lambda\phi_{\mathrm{ns}\mathrm{g}}^{\circ}.$ (S32)
Thus, per the Gibbs phase rule, only one unknown $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}$(and its leading-order finite-droplet correction, $\lambda_{\mathrm{ns}\mathrm{g}}$) remain to be determined. For brevity, since we will find $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}\ll 1$, we set $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\ll e^{-\chi}$ to be simply $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}=0$.
Lastly, we confirm that the FH surface tension (Eq. (S13)) does not change due to “labeling” inert (nonsticky) species. First, Eq. (S25b) simplifies for the FH free energy as $\phi_{\mathrm{ns}}(x)/[1-\phi_{\mathrm{s}}(x)-\phi_{\mathrm{ns}}(x)]=\phi_{\mathrm{ns}\mathrm{g}}^{\circ}/(1-\phi_{\mathrm{s}\mathrm{g}}^{\circ}-\phi_{\mathrm{ns}\mathrm{g}}^{\circ})$, which is the position-dependent analogue of the inert proteome/nonsticky protein equivalence statement, Eq. (S30). Inserting it in Eq. (S27) with the FH free energy (Eq. (S28)) results in Eq. (S13) with the binary-fluid’s FH free energy (Eq. (S14)). Thus, indeed the surface tension is ignorant of inert-protein labels, so $\sigma$ is still given by Eq. (S18).
SII From volume-fraction field dynamics to droplet growth law
In this section, we detail the full derivation of the droplet growth laws in the presence of reactions. Particularly, we elaborate further on our approximations, list the phenomena that might occur in the liquid phase, derive the flux-matching equation (Eq. (21)) of the main text, extract $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ from it in the different limits of droplet size $R$, and from it the radius-growth laws. Throughout, we refer to the results of the FH free energy to support general intuition.
SII.1 Reaction-diffusion approach
Here, we construct the reaction-diffusion equations of the main text (Eq. (13)). We denote with vectors and matrices
$\bar{z}=\left(\begin{array}[]{c}z_{\mathrm{s}}\\ z_{\mathrm{ns}}\end{array}\right),\qquad\bar{\bar{Z}}=\left(\begin{array}[]{cc}Z_{\mathrm{s}\mathrm{s}}&Z_{\mathrm{s}\mathrm{ns}}\\ Z_{\mathrm{ns}\mathrm{s}}&Z_{\mathrm{ns}\mathrm{ns}}\end{array}\right),$ (S33)
to spare cumbersome notation indicating a summation over particle types.
First, following linear-irreversible thermodynamics, the volume-fraction fields $\bar{\phi}(\mathbf{r},t)$ respond linearly to chemical potential gradients $\delta F/\delta[\varrho\bar{\phi}(\mathbf{r},t)]$ via a (normalized) mobility matrix denoted $\varrho k_{\mathrm{B}}T\bar{\bar{\Gamma}}(\bar{\phi}(\mathbf{r},t))$. Thus, the diffusive part of the dynamics is given by
$\frac{\partial\bar{\phi}(\mathbf{r},t)}{\partial t}=-\boldsymbol{\nabla}\cdot\bar{\mathbf{J}}(\mathbf{r},t),\qquad\bar{\mathbf{J}}(\mathbf{r},t)=\bar{\bar{\Gamma}}(\bar{\phi}(\mathbf{r},t))\boldsymbol{\nabla}\left[\frac{\partial f}{\partial\bar{\phi}}\Biggr{|}_{\bar{\phi}(\mathbf{r},t)}-\bar{\bar{\kappa}}\nabla^{2}\bar{\phi}(\mathbf{r},t)\right],$ (S34)
where $\kappa_{ij}=\delta_{i\mathrm{s}}\delta_{j\mathrm{s}}\kappa_{\mathrm{s}}$. For a single-component system, this reproduces Eq. (2) of the main text.
Second, we suppose that a collection of $M_{\mathrm{R}}$ reactions take place, whose rates as a function of protein concentrations are $\{S_{m}(\bar{\phi})\}_{m=1,\ldots,M_{\mathrm{R}}}$. Denoting the set of stoichiometric constants for each species in the $m$th reaction as $\bar{\nu}_{m}$, we find the reactions’ contribution to the volume-fraction dynamics,
$\frac{\partial\bar{\phi}(\mathbf{r},t)}{\partial t}=\sum_{m=1}^{M_{\mathrm{R}}}\bar{\nu}_{m}S_{m}(\bar{\phi}(\mathbf{r},t)).$ (S35)
We stress that the reaction mechanism must conserve material. Namely, if $\mathrm{s}$ and $\mathrm{ns}$ react with some stoichiometric constants, they are bound to satisfy the same ratio among constants across all reactions. Otherwise, it would seem as if material is lost or gained spontaneously. If material is not conserved, this means that the two-component description is insufficient, and underlying ignored components would have to be considered. We assume that this is not the case, and the system is properly described as a two-component one. Hence, necessarily, we can rearrange Eq. (S35) as
$\sum_{m=1}^{M_{\mathrm{R}}}\bar{\nu}_{m}S_{m}(\bar{\phi}(\mathbf{r},t))\equiv\bar{\upsilon}s(\bar{\phi}(\mathbf{r},t)),$ (S36)
where $s(\bar{\phi})$ is a single unified reaction rate corresponding to some choice of stoichiometric coefficients $\bar{\upsilon}$ satisfying the abovementioned material balance.
Upon combining both reactions and diffusion, we find Eq. (13) of the main text,
$\frac{\partial\bar{\phi}(\mathbf{r},t)}{\partial t}=\boldsymbol{\nabla}\cdot\Biggl{\{}\bar{\bar{\Gamma}}(\bar{\phi}(\mathbf{r},t))\boldsymbol{\nabla}\Biggl{[}\frac{\partial f}{\partial\bar{\phi}}\Biggr{|}_{\bar{\phi}(\mathbf{r},t)}-\bar{\bar{\kappa}}\nabla^{2}\bar{\phi}(\mathbf{r},t)\Biggr{]}\Biggr{\}}+\bar{\upsilon}s(\bar{\phi}(\mathbf{r},t)).$ (S37)
In the main text, we assume a one-to-one stoichiometric ratio between $\mathrm{s}$ and $\mathrm{ns}$ for the interconversion reactions. Here, we proceed with a general ratio.
Since $\mathrm{s}$ is created from $\mathrm{ns}$ and vice versa, $\upsilon_{\mathrm{s}}$ and $\upsilon_{\mathrm{ns}}$ must be of opposite signs. Without loss of generality (up to rescaling of $s(\bar{\phi})$), we choose $\upsilon_{\mathrm{s}}<0$ and $\upsilon_{\mathrm{ns}}=1$; $|\upsilon_{\mathrm{s}}|$ thus determines the stoichiometric ratio. (In the main text, $\upsilon_{\mathrm{s}}=-1$.) For later convenience, we also define $\bar{\upsilon}_{\perp}$ as the orthogonal vector to $\bar{\upsilon}$, so $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\upsilon}=0$. Again without loss of generality, we choose it as $\bar{\upsilon}_{\perp}=(1,|\upsilon_{\mathrm{s}}|)^{\mathrm{T}}$ (whereas $\bar{\upsilon}=(-|\upsilon_{\mathrm{s}}|,1)^{\mathrm{T}}$).
Multiplying Eq. (S37) by $\bar{\upsilon}_{\perp}^{\mathrm{T}}$, we identify a conserved field $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\phi}(\mathbf{r},t)$, which is ignorant of the reactions. Next, we suppose that the nonsticky proteins are formed out of an initial sticky-only solution whose saturation was $\varphi^{\infty}$ (see main text). Thus, with the above choice for $\bar{\upsilon}_{\perp}$, the saturation is conserved via $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\varphi}^{\infty}=\varphi^{\infty}$. The two saturations $\bar{\varphi}^{\infty}=(\varphi^{\infty}_{\mathrm{s}},\varphi^{\infty}_{\mathrm{ns}})^{\mathrm{T}}$ far away from a droplet are determined from the reaction equilibrium, $s(\bar{\varphi}^{\infty})=0$.
SII.2 Interim simplifications
To make analytical progress from Eq. (S37), we made a number of standard simplifications in the main text [1], valid at a late stage of coarsening. We quickly recap them: First, the droplets are sparse (the interdroplet distances are much greater than their radii), so we may treat a single phase-separated droplet as if it is in isolation, diffusively exchanging material with a slightly-supersaturated dilute phase, $0<\varphi^{\infty}_{\mathrm{s}}-\phi_{\mathrm{s}}^{\circ}\ll 1$. Owing to surface tension, the droplets are spherical, with radius $R(t)$ much larger than the interface width $\xi$ (defined in Sec. SI.3). With spherical symmetry, we split the volume fraction fields to droplet interior (liquid) and exterior (gas), $\bar{\phi}_{\mathrm{l}}(r,t)=\bar{\phi}(|\mathbf{r}-\mathbf{r}_{\mathrm{CM}}|<R(t),t)$ and $\bar{\phi}_{\mathrm{g}}(r,t)=\bar{\phi}(|\mathbf{r}-\mathbf{r}_{\mathrm{CM}}|>R(t),t)$, respectively, where $\mathbf{r}_{\mathrm{CM}}$ is the droplet-center coordinate. Assuming the reactions are slow enough not to interfere with surface thermodynamics [39], the two volume fractions are always equilibrated at the interface, $\bar{\phi}_{\alpha}(R(t),t)=\bar{\phi}_{\alpha}^{\mathrm{L}}(R(t))$ ($\alpha=\mathrm{l},\mathrm{g}$). Lastly, the droplets grow quasistatically, whereby the volume fraction fields $\{\bar{\phi}_{\alpha}(r;R)\}$ are found for a quasi-steady state of a fixed radius $R$, and then droplet growth rate is inferred from the resulting steady diffusive flux.
In what follows, we present a more thorough treatment of the last approximation we have made in the main text — linearizing Eq. (S37) deep in each bulk. We now present the general scheme, and in subsequent sections we solve the resulting equations, discuss their validity in each phase, and make adjustments accordingly.
The linearization is based on the assumption that the volume fractions fields $\{\bar{\phi}_{\alpha}(r;R)\}$ remain close to their coexistence values, $\{\bar{\phi}_{\alpha}^{\mathrm{L}}(R)\}$, in each phase $\alpha=\mathrm{l},\mathrm{g}$. This allows us to write
$\bar{\phi}_{\alpha}(r;R)\equiv\bar{\phi}_{\alpha}^{\mathrm{L}}(R)+\delta\bar{\phi}_{\alpha}(r;R).$ (S38)
For brevity, we will omit the reminder of the parametric dependence on $R$ henceforth. Note that, in accord with Sec. SI.4, the volume fractions $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}$ ($\simeq 1$), $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$ ($\ll 1$), and $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}$ ($\to 0$) are known explicitly, whereas $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is still unknown. As mentioned there and in Sec. III A of the main text, $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ will be determined by equating the two equations for $\mathrm{d}R/\mathrm{d}t$, obtained from the sticky and nonsticky fluxes (Eqs. (5) and (12) of the main text, and Sec. SII.5 below). To obtain the equation for $\mathrm{d}R/\mathrm{d}t$, we regard $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ as a known parameter, and express the flux $J_{\mathrm{ns}}$ in term of $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. The linearization of Eq. (S38) is thus an ansatz for $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ — that $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\simeq\varphi_{\mathrm{ns}}^{\infty}$ — which we make to be able to compute the flux $J_{\mathrm{ns}}$ in terms of $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. Once all calculations are complete and $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is found, we confirm that indeed the resulting $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is close to $\varphi_{\mathrm{ns}}^{\infty}$.
Since gradients of $\delta\bar{\phi}_{\alpha}(r)$ are over large distances of order the radius and the reaction-diffusion lengthscale, we ignore the short-distance surface tension term $\kappa_{\mathrm{s}}\nabla^{2}\bar{\phi}_{\alpha}$ and obtain the quasistatic linearized equations for $\delta\bar{\phi}_{\alpha}$,
$\frac{\bar{\bar{D}}_{\alpha}}{r^{2}}\frac{\partial}{\partial r}\left[r^{2}\frac{\partial\delta\bar{\phi}_{\alpha}(r)}{\partial r}\right]=-\bar{\upsilon}\left[s_{\alpha}+\bar{k}_{\alpha}^{\mathrm{T}}\delta\bar{\phi}_{\alpha}(r)\right],$ (S39)
where we have defined an effective diffusion matrix, zeroth-, and first-order reaction rates, all evaluated at the coexistence volume fractions of phase $\alpha$,
$\bar{\bar{D}}_{\alpha}$ $=$ $\bar{\bar{\Gamma}}(\bar{\phi}_{\alpha}^{\mathrm{L}})\frac{\partial^{2}f}{\partial\bar{\phi}\partial\bar{\phi}}\biggr{|}_{\bar{\phi}_{\alpha}^{\mathrm{L}}},$ (S40a)
$s_{\alpha}$ $=$ $s(\bar{\phi}_{\alpha}^{\mathrm{L}}),$ (S40b)
$\bar{k}_{\alpha}$ $=$ $\left.\frac{\partial s}{\partial\bar{\phi}}\right|_{\bar{\phi}_{\alpha}^{\mathrm{L}}}.$ (S40c)
To gain intuition about these quantities, we will occasionally recall $f(\bar{\phi})$ and $\{\bar{\phi}_{\alpha}^{\mathrm{L}}\}$ from the FH free energy (Sec. SI.4), and consider the simplest mobility matrix [71, 72, 68],
$\bar{\bar{\Gamma}}(\bar{\phi})=\left(\begin{array}[]{cc}\Gamma\phi_{\mathrm{s}}&0\\ 0&\Gamma\phi_{\mathrm{ns}}\end{array}\right).$ (S41)
A common mobility constant $\Gamma$ is taken for both the sticky and nonsticky particles since they are structurally similar, and hence the viscous resistance of the host (inert) fluid to their diffusion should be comparable.
We decouple Eqs. (S39) by defining the fields
$\left(\begin{array}[]{c}\delta\Phi_{\alpha}(r)\\ \delta\Psi_{\alpha}(r)\end{array}\right)\equiv\left(\begin{array}[]{c}\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\bar{D}}_{\alpha}\\ \bar{k}_{\alpha}^{\mathrm{T}}\end{array}\right)\delta\bar{\phi}_{\alpha}(r),$ (S42)
They obey
$\frac{1}{r^{2}}\frac{\partial}{\partial r}\left[r^{2}\frac{\partial\delta\Phi_{\alpha}(r)}{\partial r}\right]$ $=$ $0,$ (S43a)
$\frac{1}{r^{2}}\frac{\partial}{\partial r}\left[r^{2}\frac{\partial\delta\Psi_{\alpha}(r)}{\partial r}\right]$ $=$ $\frac{s_{\alpha}+\delta\Psi_{\alpha}(r)}{\ell_{\alpha}^{2}},$ (S43b)
where we defined the reaction-diffusion lengthscale in each phase,
$\frac{1}{\ell_{\alpha}^{2}}\equiv-\bar{k}_{\alpha}^{\mathrm{T}}\bar{\bar{D}}_{\alpha}^{-1}\bar{\upsilon}.$ (S44)
In the coming sections, we solve Eqs. (S43) and discuss the choice of sign in Eq. (S44).
SII.3 Gas phase
Since the gas is dilute, we claimed in the main text that $D_{ij\mathrm{g}}=\delta_{ij}D_{\mathrm{g}}$, where $D_{\mathrm{g}}$ is an effective single-particle diffusion constant. This is the case since both the hydrodynamic and thermodynamic couplings are weak as the sticky molecules rarely encounter each other. This is further supported by inserting the FH free energy Eq. (S28) and the mobility Eq. (S41) in Eq. (S40a),
$\bar{\bar{D}}_{\mathrm{g}}=\Gamma\left(\begin{array}[]{cc}1-(2\chi-1)\left(e^{-\chi}+\frac{\lambda}{R}\right)&e^{-\chi}+\frac{\lambda}{R}\\ \phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}&1+\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\end{array}\right),$ (S45)
where we have kept leading-order corrections in small $e^{-\chi}\ll 1$, $\lambda/R\ll 1$, and $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\ll 1$ for demonstration. Indeed, the leading order is a diagonal diffusivity matrix, $D_{ij\mathrm{g}}=\delta_{ij}D_{\mathrm{g}}$, with $D_{\mathrm{g}}=\Gamma$.
We proceed to investigate the reactions, $\partial\bar{\phi}(t)/\partial t=\bar{\upsilon}s(\bar{\phi}(t))$. Once a total amount of material is imposed (e.g., in our case, $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\varphi}^{\infty}=\varphi^{\infty}$), chemical equilibrium is determined by the condition $s(\bar{\varphi}^{\infty})=0$. For it to be a stable equilibrium, we require that the Jacobian, $\bar{\upsilon}(\partial s/\partial\bar{\phi})^{\mathrm{T}}|_{\bar{\varphi}^{\infty}}\simeq\bar{\upsilon}\bar{k}_{\mathrm{g}}^{\mathrm{T}}$ does not have any positive eigenvalues; otherwise, if the volume fractions deviate from $\bar{\varphi}^{\infty}$ (which is indeed the case around a surface, where the volume fractions are slightly different, $\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$), the entire dilute phase would destabilize due to the mere presence of reactions. We make the reasonable assumption that the reactions in isolation are indeed stable. The Jacobian’s eigenvalues are 00 (corresponding to steady state) and $\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}$, so we require $\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}<0$ (corresponding to a decaying mode). Since $\bar{\bar{D}}_{\mathrm{g}}$ is to leading order diagonal, we confirm that $\ell^{2}_{\mathrm{g}}\simeq-D_{\mathrm{g}}\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}>0$ (as $D_{\mathrm{g}}>0$), so the choice of sign in Eq. (S44) is justified.
SII.3.1 Solution of the linearized equations
With the above preparations, we proceed to solve Eqs. (S43). We recall the boundary conditions for the gas phase — there is no deviation from coexistence at the interface, and far from the droplet the volume fractions account for the overall saturation $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\varphi}^{\infty}=\varphi^{\infty}$ and also equilibrate the reactions ($s_{\mathrm{g}}(\bar{\varphi}^{\infty})=s_{\mathrm{g}}+\bar{k}_{\mathrm{g}}^{\mathrm{T}}\delta\bar{\phi}_{\mathrm{g}}(\infty)=0$),
$\delta\Phi_{\mathrm{g}}(R)$ $=$ $0,\qquad\delta\Phi_{\mathrm{g}}(\infty)=\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\bar{D}}_{\mathrm{g}}\delta\bar{\phi}_{\mathrm{g}}(\infty),$ (S46a)
$\delta\Psi_{\mathrm{g}}(R)$ $=$ $0,\qquad\delta\Psi_{\mathrm{g}}(\infty)=\bar{k}_{\mathrm{g}}^{\mathrm{T}}\delta\bar{\phi}_{\mathrm{g}}(\infty)=-s_{\mathrm{g}}.$ (S46b)
The solutions to Eqs. (S43) read
$\delta\Phi_{\mathrm{g}}(r)$ $=$ $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\bar{D}}_{\mathrm{g}}\delta\bar{\phi}_{\mathrm{g}}(\infty)\left(1-\frac{R}{r}\right),$ (S47a)
$\delta\Psi_{\mathrm{g}}(r)$ $=$ $\bar{k}_{\mathrm{g}}^{\mathrm{T}}\delta\bar{\phi}_{\mathrm{g}}(\infty)\left[1-\frac{R}{r}e^{-(r-R)/\ell_{\mathrm{g}}}\right].$ (S47b)
Indeed, if the deviations $\delta\phi_{i\mathrm{g}}(\infty)=\varphi_{i}^{\infty}-\phi_{i\mathrm{g}}^{\mathrm{L}}$ are small, the profiles $\delta\Phi_{\mathrm{g}}(r)$ and $\delta\Psi_{\mathrm{g}}(r)$ are also small everywhere in the dilute phase, justifying the linearization.
We prepare the mathematical property
$\bar{\bar{D}}_{\alpha}\left(\begin{array}[]{c}\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\bar{D}}_{\alpha}\\ \bar{k}_{\alpha}^{\mathrm{T}}\end{array}\right)^{-1}\left(\begin{array}[]{c}0\\ 1\end{array}\right)=-\ell_{\alpha}^{2}\bar{\upsilon},$ (S48)
where we have kept in mind $\bar{\upsilon}=(-|\upsilon_{\mathrm{s}}|,1)^{\mathrm{T}}$, $\bar{\upsilon}_{\perp}=(1,|\upsilon_{\mathrm{s}}|)^{\mathrm{T}}$, and the definition of $\ell_{\alpha}^{2}$ (Eq. (S44)). Combining it with the definitions in Eq. (S42), we find the fluxes at the interface
$\bar{J}_{\mathrm{g}}(R)=-\bar{\bar{D}}_{\mathrm{g}}\left.\frac{\mathrm{d}\bar{\phi}_{g}}{\mathrm{d}r}\right|_{R}=-\bar{\bar{D}}_{\mathrm{g}}\left(\begin{array}[]{c}\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\bar{D}}_{\mathrm{g}}\\ \bar{k}_{\mathrm{g}}^{\mathrm{T}}\end{array}\right)^{-1}\left.\frac{\mathrm{d}}{\mathrm{d}r}\left(\begin{array}[]{c}\delta\Phi_{\mathrm{g}}\\ \delta\Psi_{\mathrm{g}}\end{array}\right)\right|_{R}=-\frac{D_{\mathrm{g}}}{R}(\bar{\varphi}^{\infty}-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}})-\bar{\upsilon}s_{\mathrm{g}}\ell_{\mathrm{g}}.$ (S49)
The first term corresponds to the passive diffusive influx of material from a supersaturated environment, where the total amount of material joining the sphere is proportional to the radius, $R^{2}J_{i\mathrm{g}}(R)\sim D_{\mathrm{g}}R$ ($i=\mathrm{s},\mathrm{ns}$). The second term is the additional flux due to reactions in the dilute phase; since reactions equilibrate past distance $\ell_{\mathrm{g}}$, the amount of material joining due to reactions with rate $s_{\mathrm{g}}$ is proportional to the shell volume surrounding the interface, $R^{2}\bar{J}_{\mathrm{g}}(R)\sim s_{\mathrm{g}}R^{2}\ell_{\mathrm{g}}$.
SII.3.2 Simpler main-text linearization
The results of the previous section stem from a general linearization procedure, and they will be used in Sec. SII.5 below. Owing to the gas phase being dilute, the results of the main text were based on a slightly simpler procedure as follows: The volume fraction fields $\bar{\phi}_{\mathrm{g}}(r)$ vary between (the small) $\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$ and (the small) $\bar{\varphi}^{\infty}$, so they are small for all $r$. Thus, we expand the reaction rates around zero volume fraction (as opposed to around $\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$ as above), $s(\bar{\phi})\simeq s(\bar{0})+(\partial s/\partial\bar{\phi})^{\mathrm{T}}|_{\bar{0}}\bar{\phi}$. There are no reactions in the absence of material, so $s(\bar{0})=0$. Next, we assume that there is a first-order reaction step in the interconversion mechanism, such that neither component of $(\partial s/\partial\bar{\phi})^{\mathrm{T}}|_{\bar{0}}$ is zero. Hence, we neglect higher-order terms in the above expansion. For first-order reactions to take place, the stoichiometry must be one to one, $\bar{\upsilon}=(-1,1)^{\mathrm{T}}$. Furthermore, we denote as in the main text, $(\partial s/\partial\bar{\phi})|_{\bar{0}}=(k_{\mathrm{s}\to\mathrm{ns}},-k_{\mathrm{ns}\to\mathrm{s}})^{\mathrm{T}}$ (with constant $k_{\mathrm{s}\to\mathrm{ns}},k_{\mathrm{ns}\to\mathrm{s}}>0$), where the sign is such that $\mathrm{s}$ is lost when creating $\mathrm{ns}$ and vice versa. Overall,
$s(\bar{\phi}_{\mathrm{g}}(r))\simeq k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}(r)-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}(r),$ (S50)
which was used on the right-hand side of Eq. (15) in the main text. This approximation simply states that only the first-order reactions occur with a meaningful rate, as encounters between molecules required for higher-order reactions are too rare. This need not imply necessarily that Eq. (S50) is the full mechanism; a different limit of the full mechanism might apply in the liquid phase.
In the language of the more general linearization of Sec. SII.2, the main-text linearization amounts to choosing $\bar{k}_{\mathrm{g}}=(k_{\mathrm{s}\to\mathrm{ns}},-k_{\mathrm{ns}\to\mathrm{s}})^{\mathrm{T}}$ and $s_{\mathrm{g}}=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{L}$. Using the diagonal diffusivity $D_{ij\mathrm{g}}=\delta_{ij}D_{\mathrm{g}}$ and the stoichiometry $\bar{\upsilon}=(-1,1)^{\mathrm{T}}$, we obtain $1/\ell_{\mathrm{g}}^{2}=(|\bar{\upsilon}_{\mathrm{s}}|k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})/D_{\mathrm{g}}$ (which is Eq. (19) of the main text). Upon inserting these facts in Eqs. (S47), we find
$\delta\phi_{\mathrm{s}\mathrm{g}}(r)+\delta\phi_{\mathrm{ns}\mathrm{g}}(r)$ $=$ $(\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})\left(1-\frac{R}{r}\right),$ (S51a)
$k_{\mathrm{s}\to\mathrm{ns}}\delta\phi_{\mathrm{s}\mathrm{g}}(r)-k_{\mathrm{ns}\to\mathrm{s}}\delta\phi_{\mathrm{ns}\mathrm{g}}(r)$ $=$ $-(k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})\left[1-\frac{R}{r}e^{-(r-R)/\ell_{\mathrm{g}}}\right],\quad$ (S51b)
which coincide with Eqs. (17) and (18) of the main text. Explicitly,
$\phi_{\mathrm{s}\mathrm{g}}(r)$ $=$ $\varphi_{\mathrm{s}}^{\infty}-\left[\varphi_{\mathrm{s}}^{\infty}-\frac{k_{\mathrm{ns}\to\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}+\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}\right]\frac{R}{r}+\frac{k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}\frac{R}{r}e^{-(r-R)/\ell_{\mathrm{g}}},\quad$ (S52a)
$\phi_{\mathrm{ns}\mathrm{g}}(r)$ $=$ $\varphi_{\mathrm{ns}}^{\infty}-\left[\varphi_{\mathrm{ns}}^{\infty}-\frac{k_{\mathrm{s}\to\mathrm{ns}}(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}+\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}\right]\frac{R}{r}-\frac{k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}\frac{R}{r}e^{-(r-R)/\ell_{\mathrm{g}}}.\quad$ (S52b)
where $\bar{\varphi}^{\infty}=(k_{\mathrm{ns}\to\mathrm{s}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}),k_{\mathrm{s}\to\mathrm{ns}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}))^{\mathrm{T}}$. As mentioned earlier, the first two terms are the standard diffusive profile around a droplet, whereas the last terms oppose departure from the chemical equilibrium, increasing the diffusive flux out of small droplets and into large droplets. To see that explicitly, the fluxes from Eq. (S49) simplify as
$J_{\mathrm{s}\mathrm{g}}(R)$ $=$ $-\frac{D_{\mathrm{g}}}{R}(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}})+\frac{D_{\mathrm{g}}}{\ell}\frac{k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}},$ (S53a)
$J_{\mathrm{ns}\mathrm{g}}(R)$ $=$ $-\frac{D_{\mathrm{g}}}{R}(\varphi_{\mathrm{ns}}^{\infty}-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})-\frac{D_{\mathrm{g}}}{\ell}\frac{k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}.$ (S53b)
which coincide with Eq. (20) of the main text.
In the derivations to follow, we use the more general linearized result of Sec. SII.3.1. It accounts for cases where a higher-order reaction dominates, namely, if $(\partial s/\partial\bar{\phi})|_{\bar{0}}=\bar{0}$. For example, $s(\bar{\phi})=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}}^{2}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}}^{2}$ with $\bar{\upsilon}=(-1,1)$, which describes a reaction mechanism where a collision between two particles of the same form is needed to form two of the other form. In that case, one should have continued the above series to $s(\bar{\phi})\simeq\bar{\phi}^{\mathrm{T}}(\partial^{2}s/\partial\bar{\phi}\partial\bar{\phi}^{\mathrm{T}})|_{\bar{0}}\bar{\phi}$, such that $\bar{k}_{\mathrm{g}}^{\mathrm{T}}=2\bar{\phi}_{\mathrm{g}}^{\mathrm{L}\mathrm{T}}(\partial^{2}s/\partial\bar{\phi}\partial\bar{\phi}^{\mathrm{T}})|_{\bar{0}}$ and $s_{\mathrm{g}}=\bar{\phi}_{\mathrm{g}}^{\mathrm{L}\mathrm{T}}(\partial^{2}s/\partial\bar{\phi}\partial\bar{\phi}^{\mathrm{T}})|_{\bar{0}}\bar{\phi}_{\mathrm{g}}^{L}=(1/2)\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$ (as opposed to the above, effectively first-order treatment where $s_{\mathrm{g}}=\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$). In total, the general linearization of Sec. SII.2 accounts for non-first-order reaction mechanisms, which is a slightly-more general formalism than the one in the main text.
SII.4 Liquid phase
In the liquid phase, the diffusion matrix is not diagonal. For example, for the FH free energy,
$\bar{\bar{D}}_{\mathrm{l}}=\Gamma\left(\begin{array}[]{cc}e^{\chi}(1-2\chi e^{-\chi}+\phi_{\mathrm{ns}\mathrm{g}}^{\circ})&e^{\chi}(1-e^{-\chi}+\phi_{\mathrm{ns}\mathrm{g}}^{\circ})\\ 0&1\end{array}\right),$ (S54)
where we have kept up to first-order corrections in small $e^{-\chi}\ll 1$ and $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}\ll 1$ and we omitted Laplace-pressure corrections for simplicity. This matrix incorporates both the mobility and the free-energy response to fluctuations in the volume fraction fields. Since the liquid phase is dense in sticky protein, the sticky proteins quickly ($\sim\Gamma e^{\chi}$) rearrange in response to a perturbation in either the sticky or nonsticky volume fraction. (At the same time, the nonsticky protein is inert, so $D_{\mathrm{ns}\mathrm{s}\mathrm{l}}\sim 0$ and $D_{\mathrm{ns}\mathrm{ns}\mathrm{l}}\sim 1$ as if it is still an ideal gas.) This dominant order-$e^{\chi}$ term arises from the second-order derivative of the entropy contribution to the FH free energy, as the sticky volume fraction is very close to $1$ (within $\sim e^{-\chi}$) and so approaching it further is entropically costly. Interestingly, since the sticky proteins attract each other, a sticky fluctuation is slightly more stable than a nonsticky fluctuation (cf. $D_{\mathrm{s}\mathrm{s}\mathrm{l}}-D_{\mathrm{s}\mathrm{ns}\mathrm{l}}=-\Gamma(2\chi-1)<0$ with $\chi\gg 1$) as the former is slightly stabilized by the energy term. This will play a role in determining $\ell_{\mathrm{l}}^{2}$ in Sec. SII.4.3. Generally, the liquid-phase analogue for the reaction-stability criterion ($\bar{k}_{\mathrm{l}}^{\mathrm{T}}\bar{\upsilon}<0$) does not impose anything on $\ell_{\mathrm{l}}^{2}=-\bar{k}_{\mathrm{l}}^{\mathrm{T}}\bar{\bar{D}}_{\mathrm{l}}^{-1}\bar{\upsilon}$ since $\bar{\bar{D}}_{\mathrm{l}}$ is not diagonal. We proceed to Secs. SII.4.1 and SII.4.2 as if the sign convention in Eq. (S44) is correct; this will not be the case in Sec. SII.4.3.
SII.4.1 Solution of the linearized equations
We recall the boundary conditions for the liquid phase — no deviation from coexistence at the interface, and simply nondiverging values at the droplet center,
$\delta\Phi_{\mathrm{l}}(R)$ $=$ $0,\qquad\delta\Phi_{\mathrm{l}}(0)<\infty,$ (S55a)
$\delta\Psi_{\mathrm{l}}(R)$ $=$ $0,\qquad\delta\Psi_{\mathrm{l}}(0)<\infty.$ (S55b)
The solutions to Eqs. (S43) read
$\delta\Phi_{\mathrm{l}}(r)$ $=$ $0,$ (S56a)
$\delta\Psi_{\mathrm{l}}(r)$ $=$ $-s_{\mathrm{l}}\left[1-\frac{R}{r}\frac{\sinh(r/\ell_{\mathrm{l}})}{\sinh(R/\ell_{\mathrm{l}})}\right].$ (S56b)
In Secs. SII.4.2 and SII.4.3, we examine the validity of the linearization approximation, and find two destabilizing effects that one must avoid in order to consider the possibility of acceleration.
Supposing momentarily that the linearization is valid, using the relation in Eq. (S48) and the definitions in Eq. (S42), we find the fluxes at the interface
$\bar{J}_{\mathrm{l}}(R)=-\bar{\bar{D}}_{\mathrm{l}}\left.\frac{\mathrm{d}\bar{\phi}_{\mathrm{l}}}{\mathrm{d}r}\right|_{R}=-\bar{\bar{D}}_{\mathrm{l}}\left(\begin{array}[]{c}\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\bar{D}}_{\mathrm{l}}\\ \bar{k}_{\mathrm{l}}^{\mathrm{T}}\end{array}\right)^{-1}\left.\frac{\mathrm{d}}{\mathrm{d}r}\left(\begin{array}[]{c}\delta\Phi_{\mathrm{l}}\\ \delta\Psi_{\mathrm{l}}\end{array}\right)\right|_{R}=\bar{\upsilon}s_{\mathrm{l}}\ell_{\mathrm{l}}\left[\coth\left(\frac{R}{\ell_{\mathrm{l}}}\right)-\frac{\ell_{\mathrm{l}}}{R}\right].$ (S57)
Since the droplet is finite sized, in contrast to the gas phase, the reactions might not equilibrate at the core. This manifests as follows: If the droplet is smaller than the inter-reaction distance of the liquid phase, $R\ll\ell_{\mathrm{l}}$, then the volume fractions are almost uniform everywhere inside the droplet, and the material is uniformly converted with rate $s_{\mathrm{l}}$ throughout the entire droplet. Thus, the flux is proportional to the volume of the droplet, $R^{2}\bar{J}_{\mathrm{l}}(R)\sim s_{\mathrm{l}}R^{3}$. Explicitly,
$\bar{J}_{\mathrm{l}}(R\ll\ell_{\mathrm{l}})\simeq\frac{1}{3}\bar{\upsilon}s_{\mathrm{l}}R.$ (S58)
In the other limit, when the droplet is sufficiently large $R\gg\ell_{\mathrm{l}}$, the particles will typically have undergone multiple interconversion reactions past a distance $\ell_{\mathrm{l}}$ into the droplet, and will have reached chemical equilibrium. Thus, net interconversions occur only in a shell of this width, $R^{2}\bar{J}_{\mathrm{l}}(R)\sim s_{\mathrm{l}}R^{2}\ell_{\mathrm{l}}$. Explicitly,
$\bar{J}_{\mathrm{l}}(R\gg\ell_{\mathrm{l}})\simeq\bar{\upsilon}s_{\mathrm{l}}\ell_{\mathrm{l}}\left(1-\frac{\ell_{\mathrm{l}}}{R}\right).$ (S59)
SII.4.2 Example: Crossing the spinodal
We consider the simplest conversion reaction mechanism, $s(\bar{\phi})=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}}$ (with $\bar{\upsilon}=(-1,1)^{\mathrm{T}}$ and $\bar{\upsilon}_{\perp}=(1,1)^{\mathrm{T}}$) with the FH free energy. In this case, $s_{\mathrm{l}}=k_{\mathrm{s}\to\mathrm{ns}}+\mathcal{O}(e^{-\chi})$, $\bar{k}_{\mathrm{l}}=(k_{\mathrm{s}\to\mathrm{ns}},-k_{\mathrm{ns}\to\mathrm{s}})^{\mathrm{T}}$, and upon inverting the nondiagonal diffusivity of Eq. (S54), $1/\ell_{\mathrm{l}}^{2}=(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})/\Gamma+\mathcal{O}(\chi e^{-\chi})$. Inserting these expressions in Eq. (S56), and inverting Eq. (S42) using the property of Eq. (S48), we find
$\bar{\phi}_{\mathrm{l}}(r)$ $=$ $\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}+k_{\mathrm{s}\to\mathrm{ns}}\ell_{\mathrm{l}}^{2}\bar{\bar{D}}_{\mathrm{l}}^{-1}\bar{\upsilon}\left[1-\frac{R}{r}\frac{\sinh(r/\ell_{\mathrm{l}})}{\sinh(R/\ell_{\mathrm{l}})}\right]$ (S62)
$=$ $\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}\left(\begin{array}[]{c}-1\\ 1\end{array}\right)\left[1-\frac{R}{r}\frac{\sinh(r/\ell_{\mathrm{l}})}{\sinh(R/\ell_{\mathrm{l}})}\right].$
In contrast to the gas phase (Sec. SII.3.2), $s_{\mathrm{l}}\sim 1\gg s_{\mathrm{g}}\sim\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}$. As a result, we see in Eq. (S62) that the deviations from the coexistence (binodal) volume fractions $\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}$ might be of order up to $k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\sim 1$. When the droplets are small ($R\ll\ell_{\mathrm{l}}$), the second term would be of order $[k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})](R^{2}-r^{2})/\ell_{\mathrm{l}}^{2}$ which is small ($R^{2}/\ell_{\mathrm{l}}^{2}\ll 1$), so the linearization approach is indeed justified. On the other hand, for a large droplet, the volume fractions at the core are modified by the above factor $k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})$. Therefore, if reactions favor the formation of nonsticky species ($k_{\mathrm{s}\to\mathrm{ns}}>k_{\mathrm{ns}\to\mathrm{s}}$, which is what is needed to obtain a meaningful acceleration), the second term in Eq. (S62) originating from the reactions becomes of order $\sim 1$. This (i) invalidates the linearization approach, and (ii) even worse, suggests that if reactions dominate over the compressibility, this might drive the volume fractions beyond the spinodal, causing the whole dense-phase to destabilize. We recall that in this situation, where $s_{\mathrm{l}}\sim\Gamma/\ell_{\mathrm{l}}^{2}\gg\Gamma(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}})/\ell_{\mathrm{l}}^{2}$, we concluded in the main text (Sec. IV A) that size control takes place [3], and hence this instability is averted. Although, no acceleration would be achieved in this case.
SII.4.3 Example: Turing instability
To potentially avoid size control, we consider another toy mechanism which suppresses reactions inside the droplet: $s(\bar{\phi})=(1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}})(k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}})$ with $\bar{\upsilon}=(-1,1)^{\mathrm{T}}$. It might describe a situation where a kinase and phosphatase do not have as much access to the dense phase as they have to the dilute phase. Instead, being a part of the cell’s proteome, they only have the available volume fraction $\phi_{\mathrm{o}}=1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}}$ inside droplets. (Of the proteome, they constitute some fraction which is absorbed into the rates $k_{\mathrm{s}\to\mathrm{ns}}$ and $k_{\mathrm{ns}\to\mathrm{s}}$.) Combining, to convert $i\to j$ with rate $k_{i\to j}$, we need an encounter between the right enzyme (available volume fraction $1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}}$) and the substrate ($\phi_{i}$). (This mechanism is indeed chemically stable for all volume fractions, $(\partial s/\partial\bar{\phi})^{\mathrm{T}}\bar{\upsilon}=-(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})(1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}})<0$, where $\phi_{\mathrm{o}}=1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}}>0$).) In the gas phase, this mechanism reverts to Eq. (S50) (since $\phi_{\mathrm{s}\mathrm{g}}+\phi_{\mathrm{ns}\mathrm{g}}\ll 1$). By construction, this mechanism suppresses the total zeroth-order reaction rate in the dense phase: To leading order in small $e^{-\chi}$, it reads, $s_{\mathrm{l}}=k_{\mathrm{s}\to\mathrm{ns}}e^{-\chi}$, i.e., sticky proteins are converted to nonsticky ones as if the whole droplet consists of only sticky proteins (nearly-$1$ volume fraction), though the enzymes only have a small available volume fraction ($e^{-\chi}$) to access the droplet. Furthermore, the linear corrections to the zeroth-order reaction rate due to fluctuations in the sticky and nonsticky volume fractions read $\bar{k}_{\mathrm{l}}=(-k_{\mathrm{s}\to\mathrm{ns}}+2k_{\mathrm{s}\to\mathrm{ns}}e^{-\chi},-k_{\mathrm{s}\to\mathrm{ns}}+(k_{\mathrm{s}\to\mathrm{ns}}-k_{\mathrm{ns}\to\mathrm{s}})e^{-\chi})^{\mathrm{T}}$. That is, $s_{\mathrm{l}}$ is an overestimate of how much sticky material is being converted to nonsticky; due to the strong exclusion, once there is a downward fluctuation in either $\mathrm{s}$ or $\mathrm{ns}$ volume fraction, the enzymes would have extra volume to enter the droplet and create back sticky particles with a rate $k_{\mathrm{s}\to\mathrm{ns}}$.
These expressions, combined with the diffusion matrix of Eq. (S48), are inserted in Eq. (S44) yield $1/\ell_{\mathrm{l}}^{2}=e^{-\chi}[(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})/\Gamma][1-2\chi k_{\mathrm{s}\to\mathrm{ns}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})]$. This expression can be understood as follows: First, the combination of very little access for interconversion enzymes into droplets ($s\sim 1-\phi_{\mathrm{s}}-\phi_{\mathrm{ns}}\ll 1$) with the rapid relaxation of fluctuations in sticky volume fraction field ($D_{\mathrm{s}i\mathrm{l}}\sim e^{\chi}$), implies that the typical distance between reactions is large, $\ell_{\mathrm{l}}\sim e^{\chi/2}$. Second, $\ell_{\mathrm{l}}^{2}$ in its current definition (Eq. (S44)) might become negative if $\chi>(1+k_{\mathrm{ns}\to\mathrm{s}}/k_{\mathrm{s}\to\mathrm{ns}})/2$, i.e., if the nonsticky protein is not converted to the sticky form fast enough. This is a result of a Turing instability, whereby diffusion destabilizes the otherwise stable reactions ($(\partial s/\partial\bar{\phi})^{\mathrm{T}}\bar{\upsilon}<0$). Specifically, as we identified in Eq. (S54), the entropic response of the sticky volume-fraction field to a fluctuation in the nonsticky one is faster than the response to a fluctuation in the sticky one (since the latter is slightly stabilized energetically, parametrized with $\chi$). Thus, unless nonsticky proteins are converted back fast enough, the sticky proteins would separate from regions where nonsticky proteins are created. Since surface thermodynamics maintain $\phi_{\mathrm{s}\mathrm{l}}(R)=\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}$ at the interface, this instability would start at the core of the droplet every distance “$\mathrm{i}\ell_{\mathrm{l}}$”. Due to the surface tension, we might expect that this would culminate in bubbles [9].
Mathematically, Turing instability manifests like so: First, we amend Eq. (S44) to $\ell_{\mathrm{T}}=\mathrm{i}\ell_{\mathrm{l}}$,
$\frac{1}{\ell_{\mathrm{T}}^{2}}\equiv\bar{k}_{\mathrm{l}}^{\mathrm{T}}\bar{\bar{D}}_{\mathrm{l}}^{-1}\bar{\upsilon}.$ (S63)
The linearized Eq. (S43b) modifies to
$\frac{1}{r^{2}}\frac{\partial}{\partial r}\left[r^{2}\frac{\partial\delta\Psi_{\mathrm{l}}(r)}{\partial r}\right]=-\frac{s_{\mathrm{l}}+\delta\Psi_{\mathrm{l}}(r)}{\ell_{\mathrm{T}}^{2}}.$ (S64)
With the boundary conditions of Eq. (S55b), the previous solution (Eq. (S56b)) becomes instead
$\delta\Psi_{\mathrm{l}}(r)=-s_{\mathrm{l}}\left[1-\frac{R}{r}\frac{\sin(r/\ell_{\mathrm{T}})}{\sin(R/\ell_{\mathrm{T}})}\right].$ (S65)
Indeed, when the droplets are larger than the Turing length, $R>\pi\ell_{\mathrm{T}}$, the instabilities start forming, and according to Eq. (S65) the volume-fraction fields found from the linear theory exhibit divergence. As in Sec. SII.4.2, this indicates that the linear theory is no longer valid, and the volume fraction fields might cross the spinodal to form “dilute pockets” inside the droplets, whose wavelength is $1/\ell_{\mathrm{T}}$.
However, for small droplets ($R\ll\ell_{\mathrm{T}}$), the patterns are yet to develop, and instead the volume fractions are roughly uniform inside the droplet, as in Sec. SII.4.1. Replacing $-\ell_{\mathrm{l}}^{2}\to\ell_{\mathrm{T}}^{2}$ in Eq. (S48), the previously-obtained flux (Eq. (S57)) becomes instead
$\bar{J}_{\mathrm{l}}(R)=\bar{\upsilon}s_{\mathrm{l}}\ell_{\mathrm{T}}\left[\frac{\ell_{\mathrm{T}}}{R}-\cot\left(\frac{R}{\ell_{\mathrm{T}}}\right)\right].$ (S66)
In the limit $R\ll\ell_{\mathrm{T}}$, we once again obtain Eq. (S58), whereby the material flux is due to predominant conversion of sticky protein into nonsticky form with rate proportional to the droplet volume, $R^{2}\bar{J}_{\mathrm{l}}(R)\sim k_{\mathrm{s}\to\mathrm{ns}}R^{3}$.
SII.4.4 Linearization-free treatment
If the reactions destabilize the liquid phase (whether due to the passing of the spinodal line or through a Turing-instability mechanism), discussing accelerated ripening becomes meaningless. Thus, we assume that neither of these effects occurs, and the free energy is able to oppose any reaction-related instabilities. Overall, we are motivated to avoid the linearization in the liquid phase.
We revert back to the quasi-static approximation (see Sec. SII.2). With spherical symmetry, our starting point is the steady-state equation for the volume-fraction field,
$0=-\frac{1}{r^{2}}\frac{\partial}{\partial r}[r^{2}\bar{J}_{\mathrm{l}}(r)]+\bar{\upsilon}s(\bar{\phi}_{\mathrm{l}}(r)),$ (S67)
where $J_{\mathrm{l}}(r)$ is the (unlinearized) steady-state flux within the dense phase for a fixed radius $R$, and the boundary conditions are given in Eqs. (S55). We multiply both sides by $r^{2}$ and integrate from 00 to $R$, which results in the general expression
$R^{2}\bar{J}_{\mathrm{l}}(R)=\bar{\upsilon}\int_{0}^{R}\mathrm{d}rr^{2}s(\bar{\phi}_{\mathrm{l}}(r)).$ (S68)
We posit that the reactions have a certain reaction-diffusion lengthscale $\ell_{\mathrm{l}}$ setting the reaction equilibration distance inside the droplet. Since we have abandoned the linearized approach, it is not given by Eq. (S44), and it is instead “phenomenological”. We determine it as follows: For $R\ll\ell_{\mathrm{l}}$, the reactions do not equilibrate at the droplet core, and so we approximate $\bar{\phi}_{\mathrm{l}}(r)\simeq\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}$ per the boundary conditions set at the interface. Thus,
$\bar{J}_{\mathrm{l}}(R\ll\ell_{\mathrm{l}})\simeq\bar{\upsilon}\frac{1}{R^{2}}\int_{0}^{R}\mathrm{d}rr^{2}s(\bar{\phi}_{\mathrm{l}}^{\mathrm{L}})=\frac{1}{3}\bar{\upsilon}s_{\mathrm{l}}R,$ (S69)
where we have naturally recovered $s_{\mathrm{l}}=s(\bar{\phi}_{\mathrm{l}}^{\mathrm{L}})$. On the other hand, for $R\gg\ell_{\mathrm{l}}$, we expect the scaling $R^{2}\bar{J}_{\mathrm{l}}(R)\sim s_{\mathrm{l}}R^{2}\ell_{\mathrm{l}}$, as the reactions equilibrate within a distance $\ell_{\mathrm{l}}$ from the interface. Thus, we define $\ell_{\mathrm{l}}$ such that the following approximation is adequate:
$s(\bar{\phi}_{\mathrm{l}}(r))\simeq\begin{cases}s(\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}),&R-\ell_{\mathrm{l}}<r<R,\\ 0,&0<r<R-\ell_{\mathrm{l}}.\end{cases}$ (S70)
(For instance, in the experiment of Ref. [15], $\ell_{\mathrm{l}}$ would be the distance that water molecules can penetrate into a hydrophobic condensate.) This results in
$\bar{J}_{\mathrm{l}}(R\ll\ell_{\mathrm{l}})\simeq\bar{\upsilon}\frac{1}{R^{2}}\int_{R-\ell_{\mathrm{l}}}^{R}\mathrm{d}rr^{2}s(\bar{\phi}_{\mathrm{l}}^{\mathrm{L}})=\frac{1}{3}\bar{\upsilon}s_{\mathrm{l}}\frac{R^{3}-(R-\ell_{\mathrm{l}})^{3}}{R^{2}}\simeq\bar{\upsilon}s_{\mathrm{l}}\ell_{\mathrm{l}}\left(1-\frac{\ell_{\mathrm{l}}}{R}\right),$ (S71)
where the second approximation relied on $R\gg\ell_{\mathrm{l}}$. With this, we obtained Eqs. (S58) and (S59) for any reaction protocol (so long as it does not destabilize the liquid phase) without linearization.
For compactness, we define
$\bar{J}_{\mathrm{l}}(R)=\bar{\upsilon}s_{\mathrm{l}}\ell_{\mathrm{g}}G\left(\frac{R}{\ell_{\mathrm{g}}},\frac{\ell_{\mathrm{l}}}{\ell_{\mathrm{g}}}\right),$ (S72)
where we defined a function that interpolates between the large- and small-droplet limits,
$G\left(\frac{R}{\ell_{\mathrm{g}}},\frac{\ell_{\mathrm{l}}}{\ell_{\mathrm{g}}}\right)=\begin{cases}\frac{1}{3}\frac{R}{\ell_{\mathrm{g}}},&R\ll\ell_{\mathrm{l}},\\ \frac{\ell_{\mathrm{l}}}{\ell_{\mathrm{g}}}\left(1-\frac{\ell_{\mathrm{l}}}{\ell_{\mathrm{g}}}\frac{\ell_{\mathrm{g}}}{R}\right),&R\gg\ell_{\mathrm{l}}.\end{cases}$ (S73)
In the main text, where we have taken for simplicity $\ell_{\mathrm{l}}=\ell_{\mathrm{g}}\equiv\ell$, Eq. (21) contains $g(x)=G(x,1)$.
SII.5 Droplet growth law
Equipped with the gas phase flux, Eq. (S49) (where the linearization is adequate since the mixture is only slightly supersaturated, $0<\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}\ll 1$), and the liquid-phase flux, Eq. (S72) (which we have obtained without linearization, though we needed to assume that the reactions do not destablize the dense phase), we proceed to compute the radius growth rate, expressed from either the sticky and nonsticky fluxes. From these equations, we will extract $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ and the sought-after $\mathrm{d}R/\mathrm{d}t$.
SII.5.1 Droplet growth in terms of particle flux
First, we derive Eqs. (5) and (12) of the main text. We integrate Eq. (S37) (assuming spherical symmetry) over the entire container,
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{0}^{\infty}\mathrm{d}rr^{2}\bar{\phi}(r,t)=\bar{\upsilon}\int_{0}^{\infty}\mathrm{d}rr^{2}s(\bar{\phi}(r,t)).$ (S74)
where the flux at the boundaries is zero in a closed system. Alternatively, we split the left-hand side into liquid and gas regions, and obtain using Eq. (S37) and the Leibniz integral rule,
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{0}^{\infty}\mathrm{d}rr^{2}\bar{\phi}(r,t)$ $=$ $\frac{\mathrm{d}}{\mathrm{d}t}\int_{0}^{R^{-}(t)}\mathrm{d}rr^{2}\bar{\phi}(r,t)+\frac{\mathrm{d}}{\mathrm{d}t}\int_{R^{+}(t)}^{\infty}\mathrm{d}rr^{2}\bar{\phi}(r,t)$ (S75)
$=$ $\frac{\mathrm{d}R(t)}{\mathrm{d}t}\bar{\phi}(R^{-}(t),t)-\int_{0}^{R^{-}(t)}\mathrm{d}rr^{2}\frac{1}{r^{2}}\frac{\partial}{\partial r}\left[r^{2}\bar{J}(r,t)\right]+\bar{\upsilon}\int_{0}^{R^{-}(t)}\mathrm{d}rr^{2}s(\bar{\phi}(r,t))$
$-\frac{\mathrm{d}R(t)}{\mathrm{d}t}\bar{\phi}(R^{+}(t),t)-\int_{R^{+}(t)}^{\infty}\mathrm{d}rr^{2}\frac{1}{r^{2}}\frac{\partial}{\partial r}\left[r^{2}\bar{J}(r,t)\right]+\bar{\upsilon}\int_{R^{+}(t)}^{\infty}\mathrm{d}rr^{2}s(\bar{\phi}(r,t))$
$=$ $\frac{\mathrm{d}R(t)}{\mathrm{d}t}[\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}(R(t))-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R(t))]-[\bar{J}_{\mathrm{l}}(R(t))-\bar{J}_{\mathrm{g}}(R(t))]+\bar{\upsilon}\int_{0}^{\infty}\mathrm{d}rr^{2}s(\bar{\phi}(r,t)).$
where in the last equality we recalled the fast-surface-thermodynamics boundary condition, $\bar{\phi}_{\alpha}(R,t)=\bar{\phi}_{\alpha}^{\mathrm{L}}(R)$ ($\alpha=\mathrm{l},\mathrm{g}$), and we brought back the radius dependencies of $\bar{\phi}_{\alpha}^{\mathrm{L}}(R)$ for clarity. Upon comparing Eqs. (S74) and (S75), we cancel out the reaction term $\bar{\upsilon}\int_{0}^{\infty}\mathrm{d}rr^{2}s(\bar{\phi}(r,t))$ and isolate $\mathrm{d}R(t)/\mathrm{d}t$
$[\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}(R(t))-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R(t))]\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\bar{J}_{\mathrm{l}}(R(t))-\bar{J}_{\mathrm{g}}(R(t)),$ (S76)
which yields Eqs. (5) and (12) of the main text.
Inserting Eqs. (S49) and (S72) in Eq. (S76), we find
$[\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}(R(t))-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R(t))]\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R(t)}[\bar{\varphi}^{\infty}-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R(t))]+\bar{\upsilon}\ell_{\mathrm{g}}\left[s_{\mathrm{g}}+s_{\mathrm{l}}G\left(\frac{R(t)}{\ell_{\mathrm{g}}},\frac{\ell_{\mathrm{l}}}{\ell_{\mathrm{g}}}\right)\right].$ (S77)
The first term on the right-hand side is the contribution from passive Ostwald ripening and the second term is the consequence of reactions occurring on either side of the interface. Upon multiplying both sides of Eq. (S77) by $\bar{\upsilon}_{\perp}^{\mathrm{T}}=(1,|\upsilon_{\mathrm{s}}|)$, we find Eq. (22) of the main text in the case of general stoichiometric ratio,
$\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R(t)}[\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R(t))-|\upsilon_{\mathrm{s}}|\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R(t))],$ (S78)
where we recalled $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\varphi}^{\infty}=\varphi^{\infty}$ and kept the leading order $\bar{\upsilon}_{\perp}^{\mathrm{T}}[\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}(R)-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R)]\simeq 1$ (as $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}\simeq 1$, $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\ll 1$, and $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\to 0$). Since $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\phi}(r)$ is a conserved field which is not directly impacted by the reactions, the growth rate in term of the conserved diffusive flux adopts the functional form of passive Ostwald ripening.
We do not know yet $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)$ in Eq. (S78). By multiplying both sides of Eq. (S77) by the row vector $(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)-\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}(R),\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}(R)-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R))$ ($\simeq(0,1)$), we find to leading order in small $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},1-\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\ll 1$,
$\frac{\ell_{\mathrm{g}}}{R}[\varphi^{\infty}_{\mathrm{ns}}-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)]=-\frac{\ell_{\mathrm{g}}^{2}}{D_{\mathrm{g}}}\left[s_{\mathrm{g}}+s_{\mathrm{l}}G\left(\frac{R}{\ell_{\mathrm{g}}},\frac{\ell_{\mathrm{l}}}{\ell_{\mathrm{g}}}\right)\right],$ (S79)
which is the generalization of Eq. (21) of the main text to non-first-order reaction mechanisms in the gas phase, non-one-to-one stoichiometric ratios, and $\ell_{\mathrm{l}}\neq\ell_{\mathrm{g}}$. To obtain Eq. (21), based on Sec. SII.3.2, we define $\bar{k}_{\mathrm{g}}=(k_{\mathrm{s}\to\mathrm{ns}},-k_{\mathrm{ns}\to\mathrm{s}})^{\mathrm{T}}$ (which are constant rates in the linearization of Sec. SII.3.2), from which $\ell_{\mathrm{g}}=[D_{\mathrm{g}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})]^{1/2}$, $s_{\mathrm{g}}=\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, and assume that $\ell_{\mathrm{l}}=\ell_{\mathrm{g}}\equiv\ell$.
More generally, since $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\to 0$ is subdominant and $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}\to 1$ is determined independently of $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, neither $s_{\mathrm{l}}$ nor $\ell_{\mathrm{l}}$ depend on $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. Therefore, only the total zeroth-order reaction rate $s_{\mathrm{g}}=s(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})$ and inter-reaction distance $\ell_{\mathrm{g}}=\{D_{\mathrm{g}}/[-\bar{k}_{\mathrm{g}}^{\mathrm{T}}(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})\bar{\upsilon}]\}^{1/2}$ might depend on $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. (In the simpler case of the main text and Sec. SII.3.2, $\ell_{\mathrm{g}}$ does not after all.) Upon determining $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ from Eq. (S79), the radius growth law can be found explicitly from Eq. (S78). We proceed to do so in the different limits of $R$ versus $\ell_{\mathrm{g}}$ and $\ell_{\mathrm{l}}$.
SII.5.2 Small droplet
We begin with the limit $R\ll\ell_{\mathrm{g}},\ell_{\mathrm{l}}$. Using Eq. (S73), Eq. (S79) becomes
$\frac{\ell_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})}{R}(\varphi^{\infty}_{\mathrm{ns}}-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})=-\frac{\ell_{\mathrm{g}}^{2}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})}{D_{\mathrm{g}}}\left[s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})+\frac{s_{\mathrm{l}}}{3}\frac{R}{\ell_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})}\right],$ (S80)
where for brevity, we again suppress the $R$-dependence of $\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R)$, but bring forth the dependence of $s_{\mathrm{g}}$ and $\ell_{\mathrm{g}}$ on $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. (They also depend on $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$, but it is already known from the free energy; see Sec. SI.) Solving Eq. (S80) iteratively up to a leading-order correction for $R/\ell_{\mathrm{g}}\ll 1$ yields
$\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}=\varphi^{\infty}_{\mathrm{ns}}+\frac{\ell_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})R}{D_{\mathrm{g}}}s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})+\frac{s_{\mathrm{l}}R^{2}}{3D_{\mathrm{g}}}\simeq\varphi^{\infty}_{\mathrm{ns}}+\frac{\ell_{\mathrm{g}}(\varphi^{\infty}_{\mathrm{ns}})R}{D_{\mathrm{g}}}s_{\mathrm{g}}(\varphi^{\infty}_{\mathrm{ns}})+\frac{s_{\mathrm{l}}R^{2}}{3D_{\mathrm{g}}}.$ (S81)
Indeed, $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\simeq\varphi_{\mathrm{ns}}^{\infty}$, thus justifying the linearization ansatz (Eq. (S38)) for the nonsticky species as well. (To recover Eq. (23) of the main text from this general expression, insert $s_{\mathrm{g}}=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, $\ell_{\mathrm{g}}=[D_{\mathrm{g}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})]^{1/2}$, and $\varphi_{\mathrm{ns}}^{\infty}=k_{\mathrm{s}\to\mathrm{ns}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})$.)
Inserting this result in Eq. (S78), we find
$\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R(t)}[\varphi^{\infty}_{\mathrm{s}}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R(t))]+|\upsilon_{\mathrm{s}}|\ell_{\mathrm{g}}(\varphi^{\infty}_{\mathrm{ns}})[-s_{\mathrm{g}}(\varphi^{\infty}_{\mathrm{ns}})]-\frac{|\upsilon_{\mathrm{s}}|s_{\mathrm{l}}R(t)}{3}.$ (S82)
(To obtain Eq. (24) of the main text, insert the above conditions along with $|\upsilon_{\mathrm{s}}|=1$.) The first term corresponds to passive Ostwald ripening, the second term is a subdominant acceleration due to reactions in the dilute phase, and the third term stems from the growth-opposing reactions in the dense droplet.
SII.5.3 Large droplet
This time we assume $\ell_{\mathrm{g}},\ell_{\mathrm{l}}\ll R$. Using Eq. (S73), Eq. (S79) becomes
$\frac{\ell_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})}{R}(\varphi^{\infty}_{\mathrm{ns}}-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})=-\frac{\ell_{\mathrm{g}}^{2}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})}{D_{\mathrm{g}}}\left[s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})+s_{\mathrm{l}}\frac{\ell_{\mathrm{l}}}{\ell_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})}\left(1-\frac{\ell_{\mathrm{l}}}{R}\right)\right],$ (S83)
Here, without further assumptions (as in Sec. SII.3.2 and in the main text) $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}=\phi_{\mathrm{ns}\mathrm{g}}^{\circ}+\lambda_{\mathrm{ns}\mathrm{g}}/R$ can only be found implicitly. To leading-order in $\ell_{\mathrm{g}}/R,\lambda_{\mathrm{ns}\mathrm{g}}/R\ll 1$, $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}$ must satisfy implicitly
$-s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})\ell_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})=s_{\mathrm{l}}\ell_{\mathrm{l}}.$ (S84)
That is, since the reactions occur with rates proportional to the droplet surface, while the droplet volume can only grow at a rate proportional to the radius due to the mass conservation constraint (Eq. (S78)), the reactions must balance each other to leading order for a large droplet, $(4\pi R^{2}\ell_{\mathrm{g}})s_{\mathrm{g}}+(4\pi R^{2}\ell_{\mathrm{l}})s_{\mathrm{l}}=0$, where the corresponding reactive shell widths around the interface are $\ell_{\alpha}$ ($\alpha=\mathrm{l},\mathrm{g}$). (In the limit considered in Sec. SII.3.2 and the main text, $s_{\mathrm{g}}=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ and $\ell_{\mathrm{g}}=\ell_{\mathrm{l}}$, Eq. (S84) rearranges as $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}=(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}+s_{\mathrm{l}}/k_{\mathrm{ns}\to\mathrm{s}}$, which is the first line of Eq. (25) of the main text.) Since, to avoid size control, $s_{\mathrm{l}}<(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})[\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{L}]$, both $\mathrm{s}_{\mathrm{l}}$ and $s_{\mathrm{g}}$ are of order $\varphi^{\infty}$, and hence the expression for $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}$ will be of order $\varphi^{\infty}$; thus, the linearization (Eq. (S38)) is justified in this limit as well.
To find the next-order, we Taylor-expand $s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})$ and $\ell_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})$ around $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}$. For brevity, we will show the result for a slightly simpler case where $\ell_{\mathrm{g}}=\mathrm{const}$. We prepare $s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})=s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})+(\partial s/\partial\phi_{\mathrm{ns}})|_{\phi_{\mathrm{ns}\mathrm{g}}^{\circ}}\lambda_{\mathrm{ns}\mathrm{g}}/R=-s_{\mathrm{l}}\ell_{\mathrm{l}}/\ell_{\mathrm{g}}+k_{\mathrm{ns}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})\lambda_{\mathrm{ns}\mathrm{g}}/R$, where we used Eq. (S84) and identified the first-order coefficient $k_{\mathrm{ns}}$ (Eq. (S40c)). Thence, inserting this expansion in Eq. (S83), $\lambda_{\mathrm{ns}\mathrm{g}}$ reads
$\frac{\lambda_{\mathrm{ns}\mathrm{g}}}{R}=\frac{D_{\mathrm{g}}}{-k_{\mathrm{ns}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})\ell_{\mathrm{g}}^{2}}\left(\varphi^{\infty}_{\mathrm{ns}}-\phi_{\mathrm{ns}\mathrm{g}}^{\circ}-\frac{s_{\mathrm{l}}\ell_{\mathrm{l}}^{2}}{D_{\mathrm{g}}}\right)\frac{\ell_{\mathrm{g}}}{R}.$ (S85)
(In the case considered in the main text, where $D_{\mathrm{g}}/\ell_{\mathrm{g}}^{2}=k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}$, $k_{\mathrm{ns}}=-k_{\mathrm{ns}\to\mathrm{s}}$, $\ell_{l}=\ell_{\mathrm{g}}=\ell$, and the above expression for $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}=(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})\phi_{\mathrm{s}\mathrm{g}}^{\circ}+s_{\mathrm{l}}/k_{\mathrm{ns}\to\mathrm{s}}$, Eq. (S85) coincides with the second row of Eq. (25) of the main text. Inserting these expressions for $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}=\phi_{\mathrm{ns}\mathrm{g}}^{\circ}+\lambda_{\mathrm{ns}\mathrm{g}}/R$ in Eq. (S78) yields Eq. (26) of the main text; without these additional assumptions allowing us to express $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ explicitly, we cannot immediately conclude that acceleration occurs.) Should $\ell_{\mathrm{g}}$ also depend on $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, one would need to expand it in small $\lambda_{\mathrm{ns}\mathrm{g}}/R$ in Eq. (S83), too.
SII.5.4 Intermediate droplet
What we mean to explore in this section, for completeness, is a situation where reactions occur everywhere throughout the droplet, even for larger droplets, but in the regime where $R\ll\ell_{\mathrm{l}}$, so that the volume fraction of the particles remains constant throughout the droplet, $\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}$. (For instance, this might be the case in biocondensates, where the enzymes might be attached to the sticky species and thus always present and active inside the dense phase.) The limit where $R\ll\ell_{\mathrm{g}}$ was explored in Sec. SII.5.2. Here, we aim to explore the intermediate regime, $\ell_{\mathrm{g}}\ll R\ll\ell_{\mathrm{l}}$. Therefore, Eq. (S80) is still valid. However to avoid size control (to be able to cross from $R\ll\ell_{\mathrm{g}}$ to $R\gg\ell_{\mathrm{g}}$), we must assume that the total reaction rate inside the droplet is much smaller than in the surrounding gas phase, $s_{\mathrm{l}}R^{3}\ll|s_{\mathrm{g}}|R^{2}\ell_{\mathrm{g}}$; otherwise the last term in Eq. (S80) will strongly drive the shrinkage of the droplet.
As in Sec. SII.5.3, $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is only found implicitly. With the assumption $s_{\mathrm{l}}\ll|s_{\mathrm{g}}|\ell_{\mathrm{g}}/R$, the criterion reads
$s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})=0;$ (S86)
cf. Eq. (S84) with $s_{\mathrm{l}}\to 0$ (and $\ell_{\mathrm{g}}\neq 0$). Since in this combination of limits, reactions only create material from the gas-phase side of the interface, this equation states that the gas-phase reactions must balance out on their own. Next, we again find $\lambda_{\mathrm{ns}\mathrm{g}}$ in the case where $\ell_{\mathrm{g}}=\mathrm{const}$,
$\frac{\lambda_{\mathrm{ns}\mathrm{g}}}{R}=\frac{D_{\mathrm{g}}}{-k_{\mathrm{ns}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})\ell_{\mathrm{g}}^{2}}\left(\varphi^{\infty}_{\mathrm{ns}}-\phi_{\mathrm{ns}\mathrm{g}}^{\circ}\right)\frac{\ell_{\mathrm{g}}}{R}+\frac{s_{\mathrm{l}}}{3[-k_{\mathrm{ns}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})]}\frac{R}{\ell_{\mathrm{g}}}.$ (S87)
Thus, it is as if there are no reactions inside the droplet (cf. Eq. (S85) with $s_{\mathrm{l}}\to 0$ as the first term in Eq. (S87)), to which we add retroactively the small correction due to the subdominant dense-phase reactions.
Explicitly, in the case considered in the main text and Sec. SII.3.2, we find
$\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}=\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\left(1+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\right)(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})\frac{\ell_{\mathrm{g}}}{R}+\frac{s_{\mathrm{l}}}{3k_{\mathrm{ns}\to\mathrm{s}}}\frac{R}{\ell_{\mathrm{g}}}.$ (S88)
Inserting this in Eq. (S78) and taking the limit of late ripening (so $(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})\ell_{\mathrm{g}}\ll\lambda$), we find
$\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\frac{D^{\mathrm{eff}}}{R}[\varphi^{\infty}_{\mathrm{s}}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R(t))]-\frac{s_{\mathrm{l}}}{3k_{\mathrm{ns}\to\mathrm{s}}}\frac{D_{\mathrm{g}}}{\ell_{\mathrm{g}}},$ (S89)
which corresponds to accelerated growth of the main text (in the absence of dense-phase reactions), until the droplet is so large that the subdominant dense-phase reactions are able to oppose growth. The following final radius is reached in this case: $R^{\#}=\ell_{\mathrm{g}}(3k_{\mathrm{ns}\to\mathrm{s}}/s_{\mathrm{l}})(1+k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})[\varphi_{\mathrm{s}}^{\infty}(R^{\#})-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R^{\#})]$, where $\varphi_{\mathrm{s}}^{\infty}(R)$ is determined from material conservation (Eq. (10) of the main text).